### v2

# Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth

Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. This poor scaling can necessitate the use of overly large models to approximate even simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces an exponential number of activation patterns to manifest. Training first on these new parameters provides an initial solution that can later be refined by updating the underlying model weights. This approach allows us to produce function approximations that are several orders of magnitude better than their randomly initialized counterparts.
Machine Learning, ICML
1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it can only deactivate a neuron or apply the identity operation, it transforms a series of matrix multiplications into a multi-linear function; each possible configuration of active and inactive neurons can produce a unique linear transformation over a region of input space. Thus, for a ReLU network, the number of activation patterns and their corresponding linear regions can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.. Hence deep architectures may outperform shallow ones.
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which uses these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Initializing with an exponential number of regions is already improbable, but our pretraining strategy forcibly retains these regions where unassisted gradient descent would otherwise make short-sighted optimizations that discard them. Thus, our pretraining strategy leaves minimal work up to unassisted gradient descent, which then does not have to discover exponentially many new regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally trained network of equal size. Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
2 Related Work
This work is primarily concerned with novel training methodology, but it also possesses a significant approximation theoretic component. Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions and then converts that result back into a matrix representation.
Figure 1: (Top to bottom) Composed triangle waves; using collections of the above function to approximate $x^{2}$; derivatives of the above approximations.
2.1 Related Work in Approximation
Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989). Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019). In finite cases, the trade-off between width and depth is often studied in assessing the capability of a network to approximate a function.
Notably, there exist functions that can be represented with a sub-exponential number of neurons in a deep architecture, yet which require an exponential number of neurons in a wide and shallow architecture. For example, Telgarsky (2015) shows that deep neural networks with ReLU activations on a one-dimensional input are able to generate symmetric triangle waves with an exponential number of linear segments (shown in Figure 1 as the ReLU network $T(x)$). This network functions as follows: each layer takes a one-dimensional input on $[0,1]$, and outputs a one-dimensional signal also on $[0,1]$. The function they produce in isolation is a single symmetric triangle. Together in a network, each layer feeds its output to the next, performing function composition. Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
The novel reparameterization strategy contributed by the present work aims to extend this idea, compelling ReLU networks to compose rapidly oscillating triangular signals. The key insight is that instead of expressing each layer by its weights, we can train the location on $(0,1)$ of a triangular function’s peak and then set that layer’s weights accordingly. Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations. We emphasize that while the (Telgarsky, 2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivatives. We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
The techniques of Yarotsky (2017) and Liang & Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable. Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit. Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper. We will argue with numerical evidence that enforcing convergence of the derivative can be a good protection against overfitting in ReLU networks.
One of the appealing properties of neural networks is that they are highly modular. The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. One possible use case is to construct a multiplication gate. Perekrestenko et al. (2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low dimensional manifolds. In light of these prior papers, the present work revisits the $x^{2}$ approximation, seeking a generalization that can avoid the limitations of these works. Replacing $x^{2}$ with a more flexible class of functions may produce a suitable building block for techniques that can bypass polynomial interpolation.
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. One other interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
2.2 Related Work in Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
3 Compositional Networks
We begin by discussing how to deliberately architect the weights of a 4 neuron wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. Throughout the paper, we refer to these as compositional networks. First we will introduce their mathematical form, and then we will discuss how to encode them as ReLU networks.
We define a triangle function as
$T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases}$
where $0\leq a_{i}\leq 1$. This produces a triangular shape with a peak at $x=a_{i}$ and both endpoints at $y=0$. Each layer in a compositional network in isolation would compute these if directly fed the input signal. Its derivatives are the piecewise linear functions:
$T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
In a compositional network, the layers feed into each other. This composes triangle functions into triangle waves:
$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$ (2)
The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep we have
$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$ (3)
where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
Figure 2: The bottom right is a network representation of a triangle function. The bottom left shows that triangle function as a hidden layer of a compositional network. The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers. A full compositional network is assembled above.
To encode these functions into a ReLU network, we begin with triangle functions. This is the subnetwork on the right in Figure 2. Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are stacked the individual triangles they form will be composed, so it can be beneficial to think about each neuron in terms of its output waveform over the entire input domain $[0,1]$. $t_{1}$ neurons are always active, so they hold complete triangle waves. $t_{2}$ neurons are deactivated for small inputs, so they hold an alternating sequence of triangles and inactive regions. This can be seen in Figure 3. Note how $t_{1}$ and $t_{2}$ use identical slopes where both are active.
Figure 3: Each colored line shows the output signal of a neuron with respect to the input to the network. Colors match the corresponding neurons in Figure 2.
Naively stacking the triangle generators from Figure 2 together would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons.
There are two other neurons in the compositional network’s layers. The accumulation neuron passes a weighted sum of all previous triangle waves through each layer. If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on the derivations in Section 3.1, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, the ratios between the coefficients are distributed amongst all the weights, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the weight decay. Therefore, a fourth neuron is configured to output a constant signal. Other neurons can then use their connection to it as a bias. This neuron will then connect to itself so that it can scale down with each layer.
3.1 Differentiability of Model Output
Because the reparameterization makes use of exponentially many linear regions, the functions it produces will have high complexity and be in danger of overfitting. To mitigate this we focus on constraining the network to have a differentiable output. While this is technically impossible for a finite network since the output will be piecewise linear; it is possible for a derivative to exist in the limit of infinite depth. We show that there is a unique choice of scaling parameters that is determined by the peaks of the composed triangle functions which can accomplish this. Enforcing a smooth solution during pretraining imparts a bias towards smoothness during the second training stage that will help the model avoid overfit solutions. The condition we derive is necessary for the existence of a derivative. In the appendix we show that it is also sufficient.
Recalling the earlier definition of the network output $F(x)$, we would like to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined on all of $[0,1]$, which can only be done if the left and right hand derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ agree.
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network.
Figure 4: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
Figure 4 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next, and valleys persist. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. On negative slopes, the input to layer $i$ is reversed, so those copies are reflected. Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition. Because the triangles are situated back to back, the slopes of $W_{i}$ on each side of a valley are proportional. New valleys are scaled by $1/(1-a_{i})$, while old ones are scaled by $1/a_{i}$.
Before we begin reasoning about $F^{\prime}(x)$ we note that it can simplify the analysis to only consider the derivative of the error function $E^{\prime}(x)$.
Lemma 3.1.
for $x\in P_{i}$, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous.
Proof.
All of $W_{n}(x)$ for $n<i$ are differentiable at $x\in P_{i}$ since $x$ will lie in the interior of a linear region of $W_{n}$. Therefore, $f^{\prime}_{i}(x)=\sum_{n=0}^{i-1}s_{n}W^{\prime}_{n}(x)$ is continuous. Since $E^{\prime}_{i}+f^{\prime}_{i}=F^{\prime}$, their difference is a continuous function, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous. ∎
Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function. The left derivative will only be different by a constant factor.
Lemma 3.2.
For all $x\in P_{i}$, $E_{+}^{\prime}(x)$ and $E_{-}^{\prime}(x)$ are proportional to
$s_{i}-\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$ (4)
Proof.
Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$. The terms in the summation (4) mostly derive from the chain rule. We will use $R$ to represent $W_{i+}^{\prime}(x)$. The first term in the sum will be $Rs_{i}$. $W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply. There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$. This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$. Since this slope is positive, every term has the opposite sign as the first. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula. Note that this same reasoning applies to the left side, the initial slope $L$ will just be different. ∎
Lemma 3.3.
If $E^{\prime}(x)$ exists, it must be equal to 0.
Proof.
Let $S$ represent Equation (4), and $R$ and $L$ be the constants of proportionality for the directional derivatives. If $E^{\prime}_{+}=E^{\prime}_{-}$, then $R_{x}S=L_{x}S$ for all $x\in P_{i}$. Since $W_{i}$ is comprised of alternating positive and negatively sloped line segments, $R_{x}$ and $L_{x}$ have opposite signs. The only way to satisfy the equation then is if $S=0$. Consequently, $E^{\prime}(x)=0$ for all $x\in P_{i}$. ∎
We can now prove our main theorem. Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.
Theorem 3.4.
$F^{\prime}(x)$ is continuous only if the scaling coefficients are selected based on $a_{i}$ according to:
$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (5)
Proof.
Rewriting Equation (4) (which is equal to 0) for layers $i$ and $i+1$ in the following way:
$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1}{a_{i+2}}\left(s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}\right)$
$s_{i+1}(1-a_{i+2})=s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}$
allows for a substitution to eliminate the infinite sum
$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}$
Collecting all the terms gives
$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}}$
which simplifies to the desired result. ∎
4 Experiments
In this section, we implement our method to learn several convex one-dimensional curves. We present several important comparisons. To demonstrate how our networks make better use of depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. Additionally, we fix a set of initialization points such that the output is differentiable in the limit, and so that the output produces linear segments exponentially with depth. From these identical initialization points, we compare the effect of pretraining versus gradient descent alone. Lastly, we experiment with training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer for 1000 epochs to ensure convergence. The data are 500 evenly spaced points on the interval $[0,1]$ for each of the curves. Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. These curves are simple, and the data is noiseless and synthetic, so we do not use a train-test split for these experiments. We instead focus on learning an interpolation of the data with a low MSE to show that effective representations can be found for these functions.
4.2 Numerical Results
Table 1: Minimum and mean ($n=30$) MSE error approximating $y=x^{3}$ and $x^{11}$.
Network Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
Default Network $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
RAAI Distribution $2.14\times 10^{-5}$ $4.40\times 10^{-5}$ $3.97\times 10^{-2}$ $4.12\times 10^{-2}$
No Pretraining $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
Differentiability Not Enforced $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
Differentiability Enforced $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
Table 2: Minimum and mean ($n=30$) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
Network Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
Default Network $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
RAAI Distribution $3.59\times 10^{-5}$ $1.09\times 10^{-5}$ $3.63\times 10^{-2}$ $2.31\times 10^{-2}$
No Pretraining $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
Differentiability Not Enforced $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
Differentiability Enforced $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
Our main results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are those that rely on purely randomized initializations. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks.
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. When the scaling factors are trained separately, it can cause networks to exhibit behavior akin to overfitting. Even in our noiseless setting, we find that it is still possible to settle into local minima where the function interpolates only a small subset of the data points and fails to accurately represent the target curve in between them. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
The last trend to observe is the poor average performance of default networks. In a typical run of these experiments, around half of the default networks collapse. RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary, they do not correspond to Figure 2
4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 6: Pretraining (right) results in better structural retention in deeper layers.
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Rather than an exponential efficiency boost, depth is actually hindering these networks. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only introduce a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Inside the Compositional Networks
In Figure 6 we compare the top performing models with and without reparameterized pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure. It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors whereas pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
5 Limitations and Extensions
Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient, our networks still lose structure eventually. This suggests that our reparameterization is limited in the functions it can represent. It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision. It is in this remaining error gap that the second stage gradient descent becomes lost. In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function. It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only suitable for approximating convex functions. The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible. We show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1. Likewise, they could be used as a base component for assembly into larger networks. These composed networks could learn non-convex functions in higher dimensions, and they would at minimum be guaranteed to represent interpolating polynomials. The drawback is that they would likely become prohibitively large. Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work. Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions boundaries become dense as depth tends towards infinity. Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously. These properties both arise in our method, and are likely important considerations when seeking theoretical extensions. Lastly, it might be important to consider non-differentiable model outputs as a way to gain additional flexibility, but additional protections would need to be put in place against overfitting.
6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of training efficient nonlinear function representations; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our reparameterization and pretraining strategy to train ReLU networks to learn various nonlinear, convex, one-dimensional functions. Though further research is required to extend this technique to general multi-dimensional nonlinear functions, the present work’s principal significance lies in demonstrating the possibility of converting theoretical ReLU-based constructions into novel training procedures. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU networks and even more general types of neural networks.
Potential Broader Impact
This paper presents work whose goal is to enable more efficient neural networks. While the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
Appendix A Appendix
A.1 Sufficiency for differentiability
We can show that in addition to being necessary for continuity of the derivative, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
Theorem A.1.
$F^{\prime}(x)$ is continuous if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$.
Proof.
We begin by considering Equation (4) for layer $i$.
$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
Recall that this equation is telling us about the size of the discontinuities in the derivative as triangle waves are added. Each time a triangle wave is added, it can be thought of as subtracting the terms on the right. We will prove our result by substituting Equation (5) into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each $s_{n}$ in terms of $s_{i}$. Equation (5) gives a recurrence relation. Converting it to an explicit representation we have:
$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (6)
When we substitute this into Equation (4), three things happen: each term is divisible by $s_{i}$ so $s_{i}$ cancels out, every factor in the product except the last cancels, and $1-a_{i+1}$ cancels. This leaves
$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (7)
We will now argue that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error. Inductively we can show:
$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (8)
Intuitively, this means as long as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$
$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
using the inductive hypothesis to make the substitution
$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded and lower bounded by exponentially decaying series. Since both series approach zero, so does the series here. ∎
This lemma shows that the derivative of the finite approximation excluding $W_{i}$ is the same as that of the infinite sum.
Lemma A.2.
For all $x\in P_{i}$:
$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (9)
Proof.
From the previous lemma we know $E^{\prime}(x)=0$. $F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$. The sum of the first $i-1$ terms are differentiable at the points $P_{i}$ since they lie between the discontinuities at $B_{i-1}$. ∎
A.2 Second Derivatives
Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a continuous second derivative. To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the left and right hand limits are not equal (unless $a_{i}=0.5$), which will imply that the continuous version of the limit for the second derivative does not exist. First we will produce the series of $\Delta x$. Let $x$ be the location of a peak at layer $i$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
Lemma A.3.
If $c<a_{i}<1-c$ for all $i$, we have $\lim_{n\rightarrow\infty}r_{n}=\lim_{n\rightarrow\infty}r_{n}=x$. Furthermore, $r_{n},l_{n}\neq x_{n}$ for any finite $i$.
Proof.
Let $K$ be a constant of proportionality for the slopes on the left and right of $x$ in layer $i$. These slopes are proportional to $1/a_{i}$ and $1/(1-a_{i})$ depending on which way the triangle encompassing $x$ is oriented ($W^{\prime}_{i-1}$ could have been positive or negative at its location), We will denote these as $1/L$ and $1/R$ and reason about them later. $x_{n}$ is a peak location of $W_{i}$, so on the left side slope is negative and the right is positive. Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})L/K$ and $r_{1}=x+(1-a_{i+1})R/K$.
On each subsequent iteration $i+n$, $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$. $x$ is a valley point so the left slope is positive and the right is negative, and $l_{n}$,$r_{n}$ are peak points. The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $C_{i+n}$ oscillates from 0 to 1 over these spans. Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$. The resulting non-recursive formulas are:
$x-l_{n}=\frac{L}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{R}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}$ (10)
The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
Next we derive the values of $\Delta y$ to complete the proof.
Theorem A.4.
A compositional network with a second differentiable output necessarily outputs $y=x^{2}$.
Proof.
The points $l_{n}$ and $r_{n}$ are all peak locations, (9) gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. Earlier we reasoned about the sizes of the discontinuities in $F^{\prime}(x)$ at $x$, since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using (8) with the initial discontinuity size set to $s_{i}(K/R)$ rather than 1. Focusing on the right hand side we get:
$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=s_{i}(K/R)\prod_{m=2}^{n}(1-a_{i+m})$
taking $\Delta y/\Delta x$ gives a series:
$\frac{K^{2}s_{i}}{R^{2}(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$. The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$. ∎
A.3 Error Decay
Lastly, we show that the error of these approximations decays exponentially.
Lemma A.5.
The ratio $s_{i+2}/s_{i}$ is at most $0.25$.
Proof.
by applying formula (5) twice, we have
$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
Lemma A.6.
The function $F(x)$ is convex.
Proof.
To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i}[2^{i}]]$, which tracks the values of the derivative at the $i$th set of valley points. All but the first and last points will have been peaks at some point in their history, so Equation (9) gives the value of those derivatives as $f^{\prime}_{i}$.
We establish an inductive invariant that the y-values in the list $Y_{i}$ remain sorted in descending order, and that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$.
Before any of $W_{i}$ are added, $f_{0}$ is a line with derivative 0, $V_{0}$ is its two endpoints. $Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore the points in $Y^{\prime}$ are in descending sorted order. The second part of the invariant is true since 0 is in between those values.
Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$. The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$. This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order.
Adding $s_{i}W_{i}$ takes $f_{i}$ to $f_{i}+1$ splitting each constant valued interval in two about the points $P_{i}$, increasing the left side, and decreasing the right side. Recalling from the derivation of (4) all terms but the first in the sum have the same sign, so the values in $Y^{\prime}_{i}$ are approached monotonically. Therefore on the left interval we have $Y^{\prime}_{i}[n]>f_{i+1}>f_{i}$ and on the right we have $f_{i}>f_{i+1}>Y^{\prime}_{i}[n+1]$. And so $f_{i+1}$ remains monotone decreasing. ∎
Theorem A.7.
The approximation error $E_{i}(x)$ decays exponentially.
Proof.
To get this result, we apply the previous two lemmas. Since $F$ is convex, it lies above any line segment connecting any two points on the curve. $W_{i}(x)=1$ for all $x\in P_{i}$, but $W_{i}(x)=0$ for points in $V_{i}$. Since the bend points are only of nonzero value once, $f_{i}(x)=F(x)$ for all points in $V_{i}$. $f_{i}$ is made of line segments and equals $F$ repeatedly, $E_{i}$ will be a series of positively valued curve segments. The derivative will still be decreasing on each of these intervals since it was just shifted by a constant, and each of these intervals will be convex itself.
Since $E^{\prime}(x)=0$ for $x\in P_{i}$, these points will be the locations of maximum error. Since they only have nonzero value in $W_{i}$, and $E$ is convex, $s_{i}W_{i}(x)=max(E)=s_{i}$. Since $s_{i}$ decay exponentially, we have our result. ∎

### v3

@@
 This poor scaling can necessitate the use of overly large models to approximate even simple functions.
-To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces an exponential number of activation patterns to manifest.
-Training first on these new parameters provides an initial solution that can later be refined by updating the underlying model weights.
-This approach allows us to produce function approximations that are several orders of magnitude better than their randomly initialized counterparts.
-Machine Learning, ICML
+To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to display a number of activation patterns exponential in depth.
+Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights.
+This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
 1 Introduction
@@
 The rectified linear unit (ReLU) has a unique interpretation in this regard.
-Since it can only deactivate a neuron or apply the identity operation, it transforms a series of matrix multiplications into a multi-linear function; each possible configuration of active and inactive neurons can produce a unique linear transformation over a region of input space.
-Thus, for a ReLU network, the number of activation patterns and their corresponding linear regions can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions..
+Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear.
+However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space.
+The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network, and can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions..
 Hence deep architectures may outperform shallow ones.
@@
 Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions.
-Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized.
+Hanin and Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized.
 They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training.
 As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize.
-Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
+Gradient descent is also prone to redundancy; Frankle and Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
 The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting.
 Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain.
-We then demonstrate a novel pretraining strategy, which uses these derived parameters before manipulating the underlying matrix weights.
+We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights.
 This allows the network to discover solutions that are more accurate and unlikely to be found otherwise.
-Initializing with an exponential number of regions is already improbable, but our pretraining strategy forcibly retains these regions where unassisted gradient descent would otherwise make short-sighted optimizations that discard them.
-Thus, our pretraining strategy leaves minimal work up to unassisted gradient descent, which then does not have to discover exponentially many new regions.
-Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally trained network of equal size.
+Our technique directly rectifies the issues raised by Hanin and Rolnick (2019).
+Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise.
+Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage.
+Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions.
+Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size.
 Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
 2 Related Work
-This work is primarily concerned with novel training methodology, but it also possesses a significant approximation theoretic component.
+This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component.
 Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section.
-The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions and then converts that result back into a matrix representation.
+The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
 Figure 1: (Top to bottom) Composed triangle waves; using collections of the above function to approximate $x^{2}$; derivatives of the above approximations.
@@
 Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019).
-In finite cases, the trade-off between width and depth is often studied in assessing the capability of a network to approximate a function.
+In finite cases, the trade-off between width and depth is often studied in assessing a network’s ability to approximate (learn) a function.
 Notably, there exist functions that can be represented with a sub-exponential number of neurons in a deep architecture, yet which require an exponential number of neurons in a wide and shallow architecture.
@@
 The function they produce in isolation is a single symmetric triangle.
-Together in a network, each layer feeds its output to the next, performing function composition.
+Together in a network, each layer inputs its output to the next, performing function composition.
 Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
@@
 Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations.
-We emphasize that while the (Telgarsky, 2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
-Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy.
+We emphasize that while the Telgarsky (2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
+Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang and Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy.
 To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1.
@@
 We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
-The techniques of Yarotsky (2017) and Liang & Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable.
+The techniques of Yarotsky (2017) and Liang and Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable.
 Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit.
 Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper.
-We will argue with numerical evidence that enforcing convergence of the derivative can be a good protection against overfitting in ReLU networks.
+We will argue with numerical evidence that enforcing convergence of the derivative can reduce the number of local minima during training.
 One of the appealing properties of neural networks is that they are highly modular.
@@
 Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present.
-One other interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks.
+One other interesting comparison of the present work is to Ivanova and Kubat (1995), which uses decision trees as a means to initialize neural networks.
 It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
@@
 This stands in sharp contrast the current standard approach, which treats neurons homogeneously.
-Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010).
+Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot and Bengio, 2010).
 They use weight values that are sampled from distributions defined by the input and output dimension of each layer.
@@
 Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020).
-Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed.
+Several papers propose solutions: Shin and Karniadakis (2020) use a data-dependent initialization, while Singh and Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed.
 We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU.
@@
 3 Compositional Networks
-We begin by discussing how to deliberately architect the weights of a 4 neuron wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$.
+We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$.
 Throughout the paper, we refer to these as compositional networks.
@@
 where $0\leq a_{i}\leq 1$.
-This produces a triangular shape with a peak at $x=a_{i}$ and both endpoints at $y=0$.
+This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints at $y=0$.
 Each layer in a compositional network in isolation would compute these if directly fed the input signal.
-Its derivatives are the piecewise linear functions:
+$T_{i}(x)$’s derivatives are the piecewise linear functions:
 $T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
@@
 $W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$ (2)
-The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep we have
+The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
 $F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$ (3)
 where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
-Figure 2: The bottom right is a network representation of a triangle function.
-The bottom left shows that triangle function as a hidden layer of a compositional network.
+Figure 2: The rightmost diagram is a network representation of a triangle function.
+The middle shows that triangle function as a hidden layer of a compositional network.
 The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers.
-A full compositional network is assembled above.
+A full compositional network is assembled on the left.
 To encode these functions into a ReLU network, we begin with triangle functions.
@@
 The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope.
-When these components are stacked the individual triangles they form will be composed, so it can be beneficial to think about each neuron in terms of its output waveform over the entire input domain $[0,1]$.
-$t_{1}$ neurons are always active, so they hold complete triangle waves.
-$t_{2}$ neurons are deactivated for small inputs, so they hold an alternating sequence of triangles and inactive regions.
-This can be seen in Figure 3.
-Note how $t_{1}$ and $t_{2}$ use identical slopes where both are active.
+When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously.
+Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves.
+In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions, which can be seen in Figure 3.
 Figure 3: Each colored line shows the output signal of a neuron with respect to the input to the network. Colors match the corresponding neurons in Figure 2.
-Naively stacking the triangle generators from Figure 2 together would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep.
+Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep.
 We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead.
-Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave does so by combining $t_{1}$ and $t_{2}$ in the same proportion.
+Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion.
 In this way, we can avoid having to use the extra intermediate neurons.
+Notice in Figure 3 how the waves output by $t_{1}$ and $t_{2}$ neurons always have identical slopes in regions where both neurons are active.
 There are two other neurons in the compositional network’s layers.
-The accumulation neuron passes a weighted sum of all previous triangle waves through each layer.
+The accumulation neuron (marked as “sum” in Figure 2) acts similarly to a skip connection, maintaining a weighted sum of all previous triangle waves through each layer.
 If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients.
-Based on the derivations in Section 3.1, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues.
-Instead, the ratios between the coefficients are distributed amongst all the weights, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer.
-A conventional bias will have no connections to prior layers, so it will be unable to adjust to the weight decay.
-Therefore, a fourth neuron is configured to output a constant signal.
-Other neurons can then use their connection to it as a bias.
-This neuron will then connect to itself so that it can scale down with each layer.
-3.1 Differentiability of Model Output
-Because the reparameterization makes use of exponentially many linear regions, the functions it produces will have high complexity and be in danger of overfitting.
-To mitigate this we focus on constraining the network to have a differentiable output.
-While this is technically impossible for a finite network since the output will be piecewise linear; it is possible for a derivative to exist in the limit of infinite depth.
-We show that there is a unique choice of scaling parameters that is determined by the peaks of the composed triangle functions which can accomplish this.
-Enforcing a smooth solution during pretraining imparts a bias towards smoothness during the second training stage that will help the model avoid overfit solutions.
-The condition we derive is necessary for the existence of a derivative.
-In the appendix we show that it is also sufficient.
-Recalling the earlier definition of the network output $F(x)$, we would like to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined on all of $[0,1]$, which can only be done if the left and right hand derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ agree.
-Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$.
-We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends.
-$B_{i}:=P_{i}\cup V_{i}$.
-$f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$.
-The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network.
-Figure 4: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
-Figure 4 highlights some important properties about composing triangle functions.
-Peaks alternate with valleys.
-Peak locations in one layer become valleys in the next, and valleys persist.
-To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$.
-On negative slopes, the input to layer $i$ is reversed, so those copies are reflected.
-Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition.
-Because the triangles are situated back to back, the slopes of $W_{i}$ on each side of a valley are proportional.
-New valleys are scaled by $1/(1-a_{i})$, while old ones are scaled by $1/a_{i}$.
-Before we begin reasoning about $F^{\prime}(x)$ we note that it can simplify the analysis to only consider the derivative of the error function $E^{\prime}(x)$.
-Lemma 3.1.
-for $x\in P_{i}$, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous.
-Proof.
-All of $W_{n}(x)$ for $n<i$ are differentiable at $x\in P_{i}$ since $x$ will lie in the interior of a linear region of $W_{n}$.
-Therefore, $f^{\prime}_{i}(x)=\sum_{n=0}^{i-1}s_{n}W^{\prime}_{n}(x)$ is continuous.
-Since $E^{\prime}_{i}+f^{\prime}_{i}=F^{\prime}$, their difference is a continuous function, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous.
-∎
-Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function. The left derivative will only be different by a constant factor.
-Lemma 3.2.
-For all $x\in P_{i}$, $E_{+}^{\prime}(x)$ and $E_{-}^{\prime}(x)$ are proportional to
-$s_{i}-\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$ (4)
-Proof.
-Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in.
-To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$.
-The terms in the summation (4) mostly derive from the chain rule.
-We will use $R$ to represent $W_{i+}^{\prime}(x)$.
-The first term in the sum will be $Rs_{i}$.
-$W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply.
-There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$.
-This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$.
-Note that the second term has the opposite sign as the first.
-For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$.
-For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$.
-Since this slope is positive, every term has the opposite sign as the first.
-Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula.
-Note that this same reasoning applies to the left side, the initial slope $L$ will just be different.
-∎
-Lemma 3.3.
-If $E^{\prime}(x)$ exists, it must be equal to 0.
-Proof.
-Let $S$ represent Equation (4), and $R$ and $L$ be the constants of proportionality for the directional derivatives.
-If $E^{\prime}_{+}=E^{\prime}_{-}$, then $R_{x}S=L_{x}S$ for all $x\in P_{i}$.
-Since $W_{i}$ is comprised of alternating positive and negatively sloped line segments, $R_{x}$ and $L_{x}$ have opposite signs.
-The only way to satisfy the equation then is if $S=0$.
-Consequently, $E^{\prime}(x)=0$ for all $x\in P_{i}$.
-∎
-We can now prove our main theorem.
-Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions.
-The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks.
-Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.
+Based on the derivations in the appendix, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues.
+Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer.
+A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values.
+Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias.
+The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.
+3.1 Differentiable Model Output
+Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged with most choices of scaling parameters.
+This can lead to the network settling into overfit local minima, in which the data points might lie along the curve, but the curve is not predictive of the areas in between due to rapid oscillation.
+Our main mathematical result addresses this issue by forcing differentiability of the network output.
+It turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
 Theorem 3.1.
-$F^{\prime}(x)$ is continuous only if the scaling coefficients are selected based on $a_{i}$ according to:
+$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
 $s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (4)
-Proof.
-Rewriting Equation (4) (which is equal to 0) for layers $i$ and $i+1$ in the following way:
-$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1}{a_{i+2}}\left(s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}\right)$
-$s_{i+1}(1-a_{i+2})=s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}$
-allows for a substitution to eliminate the infinite sum
-$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}$
-Collecting all the terms gives
-$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}}$
-which simplifies to the desired result. ∎
 4 Experiments
-In this section, we implement our method to learn several convex one-dimensional curves.
-We present several important comparisons.
-To demonstrate how our networks make better use of depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both.
-Additionally, we fix a set of initialization points such that the output is differentiable in the limit, and so that the output produces linear segments exponentially with depth.
-From these identical initialization points, we compare the effect of pretraining versus gradient descent alone.
-Lastly, we experiment with training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability.
-As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
+The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function.
+Accordingly, in this section, we implement our method to learn several convex one-dimensional curves.
+These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6.
+The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
+To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh and Sreejith (2021), and produce errors that are orders of magnitude lower than both.
+For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation (4)).
+We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent.
+We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability.
+The intent of both comparisons is see if pretraining and differentiability constraints facilitate smoother navigation of the loss landscape and avoidance of local minima.
+Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
 4.1 Experimental Setup
-All models are trained using Adam (Kingma & Ba, 2017) as the optimizer for 1000 epochs to ensure convergence.
-The data are 500 evenly spaced points on the interval $[0,1]$ for each of the curves.
+All models are trained using Adam (Kingma and Ba, 2017) as the optimizer with a learning rate of 0.0001 for 1,000 epochs to ensure convergence.
 Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output.
@@
 The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable.
+As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
 The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave.
-The curves are chosen to capture a variety of convex one dimensional functions.
+The curves are chosen to capture a variety of convex one-dimensional functions.
 To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$.
@@
 This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$.
-These curves are simple, and the data is noiseless and synthetic, so we do not use a train-test split for these experiments.
-We instead focus on learning an interpolation of the data with a low MSE to show that effective representations can be found for these functions.
+The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves.
+This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network.
+We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments.
+We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible.
+The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.
 4.2 Numerical Results
-Table 1: Minimum and mean ($n=30$) MSE error approximating $y=x^{3}$ and $x^{11}$.
-Network Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
+Table 1: Minimum and mean (30 samples) MSE error approximating $y=x^{3}$ and $x^{11}$.
+Training Type Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
 Default Network $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
 RAAI Distribution $2.14\times 10^{-5}$ $4.40\times 10^{-5}$ $3.97\times 10^{-2}$ $4.12\times 10^{-2}$
-No Pretraining $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
+Pretraining Skipped $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
 Differentiability Not Enforced $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
 Differentiability Enforced $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
-Table 2: Minimum and mean ($n=30$) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
-Network Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
+Table 2: Minimum and mean (30 samples) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
+Training Type Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
 Default Network $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
 RAAI Distribution $3.59\times 10^{-5}$ $1.09\times 10^{-5}$ $3.63\times 10^{-2}$ $2.31\times 10^{-2}$
-No Pretraining $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
+Pretraining Skipped $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
 Differentiability Not Enforced $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
 Differentiability Enforced $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
-Our main results are shown in Tables 1 and 2, wherein we observe several important trends.
+Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends.
 First, the worst performing networks are those that rely on purely randomized initializations.
@@
 This indicates that these loss landscapes are indeed the most reliable to traverse.
-When the scaling factors are trained separately, it can cause networks to exhibit behavior akin to overfitting.
-Even in our noiseless setting, we find that it is still possible to settle into local minima where the function interpolates only a small subset of the data points and fails to accurately represent the target curve in between them.
+Even in our noiseless setting, we find that it is still possible to settle into local minima where the function fits only a small subset of the data points and fails to accurately represent the target curve in between them.
 Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
@@
 RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
-Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary, they do not correspond to Figure 2
+Our second set of results is shown in Table 3 and in Figure 4.
+Here the most important impact of utilizing exponentially many linear regions is demonstrated.
+Not only can more accurate representations of training data be learned, but more linear regions allow the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks.
+This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
+Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
+Table 3: Minimum errors on unseen points from training on sparse data.
+Training Type Min $x^{3}$ Min $x^{11}$ Min $\sin(x)$ Min $\tanh(3x)$
+Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
+Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
 4.3 Gradient Descent does not Directly Optimize Efficiency
-Figure 6: Pretraining (right) results in better structural retention in deeper layers.
+Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary; they do not correspond to Figure 2.
 Figure 5 shows the interior of a default network.
@@
 The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points.
-Rather than an exponential efficiency boost, depth is actually hindering these networks.
 Examining the figure, the first two layers are wasted.
@@
 Therefore, bend-related information will be absent from its gradients.
-Gradient descent will only introduce a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values.
+Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values.
 Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
-4.4 Inside the Compositional Networks
+4.4 Effect of Pretraining on a Compositional Network
+Figure 6: Pretraining (right) results in better structural retention in deeper layers.
 In Figure 6 we compare the top performing models with and without reparameterized pretraining.
 We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure.
-It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors whereas pretraining maintains structure at greater depths.
+It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors, whereas pretraining maintains structure at greater depths.
 This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results.
@@
 5 Limitations and Extensions
-Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient, our networks still lose structure eventually.
-This suggests that our reparameterization is limited in the functions it can represent.
+Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient descent, our neurons will still begin producing non-triangular outputs eventually.
+This suggests that our reparameterization is limited in the functions it can theoretically capture.
 It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision.
-It is in this remaining error gap that the second stage gradient descent becomes lost.
+It is in this remaining error gap that the second stage, gradient descent, becomes lost.
 In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
 First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function.
-It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only suitable for approximating convex functions.
+It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only guaranteed to be suitable for approximating convex functions.
 The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible.
-We show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
+The appendix show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
 Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1.
@@
 Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work.
-Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions boundaries become dense as depth tends towards infinity.
+Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions’ boundaries become dense as depth tends towards infinity.
 Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously.
@@
 6 Concluding Remarks
-This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of training efficient nonlinear function representations; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased.
+This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased.
 Our results showed improvements of one to several orders of magnitude in using our reparameterization and pretraining strategy to train ReLU networks to learn various nonlinear, convex, one-dimensional functions.
@@
 We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU networks and even more general types of neural networks.
-Potential Broader Impact
-This paper presents work whose goal is to enable more efficient neural networks.
-While the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
-Appendix A Appendix
-A.1 Sufficiency for differentiability
-We can show that in addition to being necessary for continuity of the derivative, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
-Theorem A.1.
-$F^{\prime}(x)$ is continuous if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$.
-Proof.
-We begin by considering Equation (4) for layer $i$.
-$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
-Recall that this equation is telling us about the size of the discontinuities in the derivative as triangle waves are added.
-Each time a triangle wave is added, it can be thought of as subtracting the terms on the right.
-We will prove our result by substituting Equation (5) into this formula, and then verifying that the resulting equation is valid.
-First we would like to rewrite each $s_{n}$ in terms of $s_{i}$.
-Equation (5) gives a recurrence relation.
-Converting it to an explicit representation we have:
-$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (6)
-When we substitute this into Equation (4), three things happen: each term is divisible by $s_{i}$ so $s_{i}$ cancels out, every factor in the product except the last cancels, and $1-a_{i+1}$ cancels.
-This leaves
-$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (7)
-We will now argue that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error. Inductively we can show:
-$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (8)
-Intuitively, this means as long as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
-As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$.
-Assuming the above equation holds for all previous values of $j$
-$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
-using the inductive hypothesis to make the substitution
-$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
-Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded and lower bounded by exponentially decaying series.
-Since both series approach zero, so does the series here.
-∎
-This lemma shows that the derivative of the finite approximation excluding $W_{i}$ is the same as that of the infinite sum.
-Lemma A.2.
-For all $x\in P_{i}$:
-$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (9)
-Proof.
-From the previous lemma we know $E^{\prime}(x)=0$.
-$F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$.
-The sum of the first $i-1$ terms are differentiable at the points $P_{i}$ since they lie between the discontinuities at $B_{i-1}$.
-∎
-A.2 Second Derivatives
-Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a continuous second derivative.
-To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the left and right hand limits are not equal (unless $a_{i}=0.5$), which will imply that the continuous version of the limit for the second derivative does not exist.
-First we will produce the series of $\Delta x$.
-Let $x$ be the location of a peak at layer $i$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
-Lemma A.3.
-If $c<a_{i}<1-c$ for all $i$, we have $\lim_{n\rightarrow\infty}r_{n}=\lim_{n\rightarrow\infty}r_{n}=x$. Furthermore, $r_{n},l_{n}\neq x_{n}$ for any finite $i$.
-Proof.
-Let $K$ be a constant of proportionality for the slopes on the left and right of $x$ in layer $i$.
-These slopes are proportional to $1/a_{i}$ and $1/(1-a_{i})$ depending on which way the triangle encompassing $x$ is oriented ($W^{\prime}_{i-1}$ could have been positive or negative at its location), We will denote these as $1/L$ and $1/R$ and reason about them later.
-$x_{n}$ is a peak location of $W_{i}$, so on the left side slope is negative and the right is positive.
-Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})L/K$ and $r_{1}=x+(1-a_{i+1})R/K$.
-On each subsequent iteration $i+n$, $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$.
-$x$ is a valley point so the left slope is positive and the right is negative, and $l_{n}$,$r_{n}$ are peak points.
-The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $C_{i+n}$ oscillates from 0 to 1 over these spans.
-Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$.
-The resulting non-recursive formulas are:
-$x-l_{n}=\frac{L}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{R}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}$ (10)
-The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
-Next we derive the values of $\Delta y$ to complete the proof.
-Theorem A.4.
-A compositional network with a second differentiable output necessarily outputs $y=x^{2}$.
-Proof.
-The points $l_{n}$ and $r_{n}$ are all peak locations, (9) gives their derivative values as $f^{\prime}_{i+n}(r_{n})$.
-Earlier we reasoned about the sizes of the discontinuities in $F^{\prime}(x)$ at $x$, since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using (8) with the initial discontinuity size set to $s_{i}(K/R)$ rather than 1.
-Focusing on the right hand side we get:
-$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=s_{i}(K/R)\prod_{m=2}^{n}(1-a_{i+m})$
-taking $\Delta y/\Delta x$ gives a series:
-$\frac{K^{2}s_{i}}{R^{2}(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
-The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$.
-The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$.
-∎
-A.3 Error Decay
-Lastly, we show that the error of these approximations decays exponentially.
-Lemma A.5.
-The ratio $s_{i+2}/s_{i}$ is at most $0.25$.
-Proof.
-by applying formula (5) twice, we have
-$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
-To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
-Lemma A.6.
-The function $F(x)$ is convex.
-Proof.
-To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i}[2^{i}]]$, which tracks the values of the derivative at the $i$th set of valley points.
-All but the first and last points will have been peaks at some point in their history, so Equation (9) gives the value of those derivatives as $f^{\prime}_{i}$.
-We establish an inductive invariant that the y-values in the list $Y_{i}$ remain sorted in descending order, and that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$.
-Before any of $W_{i}$ are added, $f_{0}$ is a line with derivative 0, $V_{0}$ is its two endpoints.
-$Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore the points in $Y^{\prime}$ are in descending sorted order.
-The second part of the invariant is true since 0 is in between those values.
-Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$.
-The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$.
-This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order.
-Adding $s_{i}W_{i}$ takes $f_{i}$ to $f_{i}+1$ splitting each constant valued interval in two about the points $P_{i}$, increasing the left side, and decreasing the right side.
-Recalling from the derivation of (4) all terms but the first in the sum have the same sign, so the values in $Y^{\prime}_{i}$ are approached monotonically.
-Therefore on the left interval we have $Y^{\prime}_{i}[n]>f_{i+1}>f_{i}$ and on the right we have $f_{i}>f_{i+1}>Y^{\prime}_{i}[n+1]$.
-And so $f_{i+1}$ remains monotone decreasing.
-∎
-Theorem A.7.
-The approximation error $E_{i}(x)$ decays exponentially.
-Proof.
-To get this result, we apply the previous two lemmas.
-Since $F$ is convex, it lies above any line segment connecting any two points on the curve.
-$W_{i}(x)=1$ for all $x\in P_{i}$, but $W_{i}(x)=0$ for points in $V_{i}$.
-Since the bend points are only of nonzero value once, $f_{i}(x)=F(x)$ for all points in $V_{i}$.
-$f_{i}$ is made of line segments and equals $F$ repeatedly, $E_{i}$ will be a series of positively valued curve segments.
-The derivative will still be decreasing on each of these intervals since it was just shifted by a constant, and each of these intervals will be convex itself.
-Since $E^{\prime}(x)=0$ for $x\in P_{i}$, these points will be the locations of maximum error.
-Since they only have nonzero value in $W_{i}$, and $E$ is convex, $s_{i}W_{i}(x)=max(E)=s_{i}$.
-Since $s_{i}$ decay exponentially, we have our result.
-∎

### v4

@@
-Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth
+Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training
 Abstract
@@
 For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly.
-This poor scaling can necessitate the use of overly large models to approximate even simple functions.
-To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to display a number of activation patterns exponential in depth.
+Therefore, randomly initialized models are often unnecessarily large, even when approximating simple functions.
+To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to exhibit a number of linear regions exponential in depth.
 Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights.
 This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
+We further demonstrate how to extend our approach to multidimensional and non-convex functions, with similar benefits observed.
 1 Introduction
@@
 However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space.
-The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network, and can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions..
+The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.
+and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018).
 Hence deep architectures may outperform shallow ones.
@@
 Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions.
-Hanin and Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized.
+Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized.
 They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training.
 As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize.
-Gradient descent is also prone to redundancy; Frankle and Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
+Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
 The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting.
@@
 This allows the network to discover solutions that are more accurate and unlikely to be found otherwise.
-Our technique directly rectifies the issues raised by Hanin and Rolnick (2019).
+Our technique directly rectifies the issues raised by Hanin & Rolnick (2019).
 Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise.
@@
 Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size.
-Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
+Although the preliminary results in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
 2 Related Work
 This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component.
-Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section.
+In one dimension, where we initially focus, our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section.
 The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
@@
 Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
-The novel reparameterization strategy contributed by the present work aims to extend this idea, compelling ReLU networks to compose rapidly oscillating triangular signals.
-The key insight is that instead of expressing each layer by its weights, we can train the location on $(0,1)$ of a triangular function’s peak and then set that layer’s weights accordingly.
-Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations.
-We emphasize that while the Telgarsky (2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
-Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang and Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy.
+The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)).
+Our reparameterization strategy focuses on non-symmetric triangle waves.
+The location on $(0,1)$ of a triangular function’s peak acts as an alternative parameterization of the function, instead of its weight matrix representation.
+Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions.
+Furthermore, this parameterization is trainable.
+By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions.
+The dilated triangular waveforms produced in this manner are not particularly useful on their own.
+Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit).
+But these problems can be rectified by taking a sum over the layers of a network.
+Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves.
 To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1.
-As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivatives.
-We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
-The techniques of Yarotsky (2017) and Liang and Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable.
-Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit.
-Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper.
-We will argue with numerical evidence that enforcing convergence of the derivative can reduce the number of local minima during training.
-One of the appealing properties of neural networks is that they are highly modular.
+As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$.
+Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
 The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems.
@@
 (2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$.
-The squared terms can all be moved to one side, expressing the product as a linear combination of squared terms.
+The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms.
 They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes.
@@
 (2020) uses band limiting, and Chen et al.
-(2019) approximates low dimensional manifolds.
-In light of these prior papers, the present work revisits the $x^{2}$ approximation, seeking a generalization that can avoid the limitations of these works.
-Replacing $x^{2}$ with a more flexible class of functions may produce a suitable building block for techniques that can bypass polynomial interpolation.
+(2019) approximates low-dimensional manifolds.
+These works all make use of a fixed representation of $x^{2}$.
+If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model.
+Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon.
+In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
 Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022).
 Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present.
-One other interesting comparison of the present work is to Ivanova and Kubat (1995), which uses decision trees as a means to initialize neural networks.
+Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks.
 It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
+Finally, Park et al.
+(2021) approaches the problem of linear region maximization from an information theory perspective and uses a loss penalty rather than a reparameterization to increase the number of linear regions.
 2.2 Related Work in Initialization
@@
 This stands in sharp contrast the current standard approach, which treats neurons homogeneously.
-Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot and Bengio, 2010).
+Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010).
 They use weight values that are sampled from distributions defined by the input and output dimension of each layer.
@@
 Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020).
-Several papers propose solutions: Shin and Karniadakis (2020) use a data-dependent initialization, while Singh and Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed.
+Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed.
 We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU.
 Our approach enforces a specific network structure that does not collapse in this manner.
-3 Compositional Networks
+3 Initialization and Pretraining Construction
 We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$.
-Throughout the paper, we refer to these as compositional networks.
-First we will introduce their mathematical form, and then we will discuss how to encode them as ReLU networks.
-We define a triangle function as
+As alluded to in Section 2.1, our initialization algorithm is simply to associate a nonsymmetric triangle function with each layer, and then to randomize the location of each triangle’s peak in $(0,1)$; any choice produces $2^{d}$ regions.
+This yields a reparameterization of the ReLU network in terms of the triangle peak locations, $a_{i}$, rather than in terms of the network’s weights.
+The pretraining step of our algorithm then involves learning and updating these parameters.
+This is done by setting individual neuron weights in accordance with the procedures described in this section, and then backpropagating the gradients to update the triangle peak parameters.
+Pseudocode for our entire algorithm is provided in Appendix A.1.
+To illustrate how we set weights, we first describe the mathematical functions that arise in our analysis. Triangle functions are defined as
 $T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases}$
-where $0\leq a_{i}\leq 1$.
-This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints at $y=0$.
-Each layer in a compositional network in isolation would compute these if directly fed the input signal.
+where $10<a_{i}<$ 1.
+This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$.
+Each layer considered in isolation would compute these if directly fed the input signal.
 $T_{i}(x)$’s derivatives are the piecewise linear functions:
 $T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
-In a compositional network, the layers feed into each other. This composes triangle functions into triangle waves:
+In a deep network, the layers feed into each other, composing their respective triangle functions:
 $W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$ (2)
-The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
+These triangle waves will have $2^{i}$ linear regions, doubling with each layer.
+The output of the network in pretraining will be a weighted sum over the triangle waves formed at each layer.
+Assuming the network to be infinitely deep, we have
 $F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$ (3)
 where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
-Figure 2: The rightmost diagram is a network representation of a triangle function.
-The middle shows that triangle function as a hidden layer of a compositional network.
-The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers.
-A full compositional network is assembled on the left.
-To encode these functions into a ReLU network, we begin with triangle functions.
-This is the subnetwork on the right in Figure 2.
+Figure 2: On the right is a network representation of a triangle function.
+The middle shows that triangle function as a hidden layer of a network.
+The one-dimensional input and output of a triangle function is converted into shared weights.
+A full pretraining network is assembled on the left.
+To encode these functions into the weights of a ReLU network, we begin with triangle functions (see the right subnetwork in Figure 2).
 Its maximum output is $1$ at the peak location $a\in(0,1)$.
@@
 Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves.
-In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions, which can be seen in Figure 3.
+In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions (see Figure 3).
 Figure 3: Each colored line shows the output signal of a neuron with respect to the input to the network. Colors match the corresponding neurons in Figure 2.
@@
 In this way, we can avoid having to use the extra intermediate neurons.
-Notice in Figure 3 how the waves output by $t_{1}$ and $t_{2}$ neurons always have identical slopes in regions where both neurons are active.
-There are two other neurons in the compositional network’s layers.
-The accumulation neuron (marked as “sum” in Figure 2) acts similarly to a skip connection, maintaining a weighted sum of all previous triangle waves through each layer.
-If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients.
-Based on the derivations in the appendix, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues.
+This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
+There are two other neurons in the pretraining network’s layers, each with a specific role.
+The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer.
+This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions).
+If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients.
+Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues.
 Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer.
@@
 The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.
+We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially in order to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
 3.1 Differentiable Model Output
-Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged with most choices of scaling parameters.
-This can lead to the network settling into overfit local minima, in which the data points might lie along the curve, but the curve is not predictive of the areas in between due to rapid oscillation.
+Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged or fractal with many choices of scaling parameters.
+This can lead to the network settling into solutions that are not predictive of the areas in between training points (or possibly not even interpolating the training data) due to rapid oscillation.
 Our main mathematical result addresses this issue by forcing differentiability of the network output.
-It turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
+To achieve differentiability, it turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them:
 Theorem 3.1.
@@
 $s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (4)
-4 Experiments
+4 One-Dimensional Experiments
 The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function.
@@
 The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
-To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh and Sreejith (2021), and produce errors that are orders of magnitude lower than both.
-For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation (4)).
-We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent.
+To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both.
+For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4).
+We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3).
 We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability.
-The intent of both comparisons is see if pretraining and differentiability constraints facilitate smoother navigation of the loss landscape and avoidance of local minima.
+The intent of both comparisons is see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape.
 Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
 4.1 Experimental Setup
-All models are trained using Adam (Kingma and Ba, 2017) as the optimizer with a learning rate of 0.0001 for 1,000 epochs to ensure convergence.
+All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix for a brief analysis of different learning rates).
 Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output.
 The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type.
-The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable.
-As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
+The networks unique to this paper share a common set of starting locations using the construction in Section 3, so that the effects of each training regimen are directly comparable.
+As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we focus for the moment on one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
 The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave.
@@
 Differentiability Enforced $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
+Table 3: Minimum errors on unseen points from training on sparse data.
+Training Type Min $x^{3}$ Min $x^{11}$ Min $\sin(x)$ Min $\tanh(3x)$
+Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
+Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
 Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends.
@@
 This indicates that these loss landscapes are indeed the most reliable to traverse.
-Even in our noiseless setting, we find that it is still possible to settle into local minima where the function fits only a small subset of the data points and fails to accurately represent the target curve in between them.
-Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
+Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
 The last trend to observe is the poor average performance of default networks.
@@
 RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
+Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
 Our second set of results is shown in Table 3 and in Figure 4.
@@
 This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
-Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
-Table 3: Minimum errors on unseen points from training on sparse data.
-Training Type Min $x^{3}$ Min $x^{11}$ Min $\sin(x)$ Min $\tanh(3x)$
-Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
-Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
 4.3 Gradient Descent does not Directly Optimize Efficiency
-Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary; they do not correspond to Figure 2.
+Figure 5: Approximation produced with standard methodologies (shown before ReLU is applied). The neuron colors here are arbitrary; they do not correspond to Figure 2.
 Figure 5 shows the interior of a default network.
@@
 Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values.
-Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
-4.4 Effect of Pretraining on a Compositional Network
+Occasionally, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
+4.4 Effect of Reparameterized Initialization vs Pretraining
 Figure 6: Pretraining (right) results in better structural retention in deeper layers.
-In Figure 6 we compare the top performing models with and without reparameterized pretraining.
-We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure.
-It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors, whereas pretraining maintains structure at greater depths.
+Figure 6 compares the top performing reparameterized models for $x^{3}$ with and without pretraining.
+We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure around layer 4 or 5, devolving into noisy patterns and resulting in higher errors.
+Pretraining maintains structure at greater depths.
 This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results.
-Gradient descent greedily abandons any such structure in favor of models that can be worse in the long term.
+Gradient descent greedily abandons any such structure during its optimization in favor of models that can be worse in the long term.
 A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
-5 Limitations and Extensions
-Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient descent, our neurons will still begin producing non-triangular outputs eventually.
-This suggests that our reparameterization is limited in the functions it can theoretically capture.
-It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision.
-It is in this remaining error gap that the second stage, gradient descent, becomes lost.
-In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
-First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function.
-It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only guaranteed to be suitable for approximating convex functions.
-The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible.
-The appendix show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
-Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1.
-Likewise, they could be used as a base component for assembly into larger networks.
-These composed networks could learn non-convex functions in higher dimensions, and they would at minimum be guaranteed to represent interpolating polynomials.
-The drawback is that they would likely become prohibitively large.
-Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work.
-Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions’ boundaries become dense as depth tends towards infinity.
-Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously.
-These properties both arise in our method, and are likely important considerations when seeking theoretical extensions.
-Lastly, it might be important to consider non-differentiable model outputs as a way to gain additional flexibility, but additional protections would need to be put in place against overfitting.
+This phenomenon highlights the need for theoretical work to derive more expressive training controls, which can give better guidance for the optimization of functions of greater dimension or complexity.
+5 Extension to Non-Convex Functions and Higher Dimensions
+Figure 7: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
+The results presented so far are limited to one-dimensional convex functions, but these are not limitations of our method.
+Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 9 in the Appendix for a schematic).
+For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs.
+Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
+The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al.
+(2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions).
+We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising (and of course, assembling more copies of our network into a deeper or wider larger network would give additional representational power).
+We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1.
+This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957; 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions.
+But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees.
+Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
+Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks.
+Rather than directly using our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect).
+Two slight modifications are made to each of the component networks to better interact with ReLU.
+First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU.
+Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
+For both $x^{3}-x$ and $r^{3}$, we substantially outperform standard initialization and training procedures.
+Especially striking is the increased number of linear regions in Figure 8.
+Visible in Figure 7 is the gradient descent-induced destructuring discussed in Section 4.4.
+This likely occurs due to the block diagonal structure of the larger weight matrices, which results in many zeros for the second pass of gradient descent to fill in unassisted.
+This could be addressed either by limiting training only to the block diagonal entries, or by deriving a more advanced pretraining method that makes use of the full matrix.
+Note that since $x^{3}-x$ is twice differentiable we can write it as $(x^{3}-x+3x^{2})-3x^{2}$, a difference of functions convex on $[-1,1]$ obtained by finding a parabola based on trying to make the derivative of $x^{3}-x$ monotone.
+Note that this is not a unique decomposition.
+Figure 8: Approximations of $z=\sqrt{x^{2}+y^{2}}^{3}$ using an $8\times 5$ regular network (left) and a union of two of our pretrained components (right).
+Losses are $1.5\times 10^{-4}$ and $3.5\times 10^{-6}$, respectively, demonstrating a nearly two orders of magnitude improvement using our techniques.
 6 Concluding Remarks
 This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased.
-Our results showed improvements of one to several orders of magnitude in using our reparameterization and pretraining strategy to train ReLU networks to learn various nonlinear, convex, one-dimensional functions.
-Though further research is required to extend this technique to general multi-dimensional nonlinear functions, the present work’s principal significance lies in demonstrating the possibility of converting theoretical ReLU-based constructions into novel training procedures.
+Our results showed improvements of one to several orders of magnitude in using our initialization and pretraining strategy to train ReLU networks to learn various nonlinear functions, including non-convex and multi-dimensional functions.
 This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks.
-We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU networks and even more general types of neural networks.
+Although we proposed one preliminary strategy for extending our network for higher-dimensional functions, this strategy has limitations, such as leaving many weights to be filled in by gradient descent (which it generally does not do effectively).
+We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture.
+We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU or even other networks.
+Appendix A Appendix
+Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks, as creating nonconvex functions this way will tend to make a lot of negative values.
+A.1 Initialization and Training Algorithm
+The initialization step of our algorithm is to generate a vector $A=[a_{0},a_{1},...a_{n}]^{T}$, where each $a_{i}$ is randomized in $(0,1)$.
+Given this $A$, the pretraining step of our algorithm sets the weights of the input ($I$), hidden ($H_{i}$, $11\leq i\leq n-$ 1), and output ($O$) layers of the network as follows:
+$I(x)=\begin{bmatrix}x\\ x\\ x\\ 0\end{bmatrix}+\begin{bmatrix}0\\ 0\\ -a_{0}\\ 1\end{bmatrix}$
+$H_{i}(x)=\begin{bmatrix}1&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times ReLU(H_{i-1}(x))$
+$O(x)=\begin{bmatrix}1&S_{n}/a_{n}&-S_{n}/(a_{n}-a_{n}^{2})&0\end{bmatrix}\times ReLU(H_{n-1}(x))$
+where $S_{i}$ can either be chosen independently, or chosen based on $A$.
+In the latter case, Equation 4 gives $S_{i}=s_{i}/s_{i-1}=(1-a_{i})a_{i+1}$.
+This assignment of $I$, $H$, and $O$ is used in each iteration of the pretraining algorithm.
+Algorithm 1 Initialization and Pretraining
+$A\leftarrow\text{Random}((0,1)^{n})$
+while $\text{Epochs}>0$ do
+[…]
+     Loss $\leftarrow(\text{Network}(x)-y)^{2}$
+[…]
+     A-Gradient $\leftarrow\text{Derivative}(\text{Network},A)$triangleright Backpropagate through weight setting
+[…]
+     $A\leftarrow A-\epsilon\times\text{Gradient}$triangleright Update A, Not the network
+end while
+The network weights can then be initialized once more based on the learned vector $A$ so that the network weights can be updated via regular gradient descent.
+In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs.
+Full pseudocode is listed in Algorithm 1.
+A.2 Derivations
+For convenience we first restate the functions defined in the main body of the paper.
+$T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases}$
+$T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$
+$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$
+$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$
+Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$.
+We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends.
+$B_{i}:=P_{i}\cup V_{i}$.
+$f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$.
+The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network.
+Although the derivative of an infinite sum of functions is not always equal to the sum of the individual functions derivatives, it is in this case, so notationally we use $F^{\prime}$ to describe the latter for brevity.
+Figure 10: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
+The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and on all of $[0,1]$.
+Fortunately, the left and right derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist, since each bend in each $W_{i}$ will use the slope of the line segment to its left or right respectively.
+The problem is that without choosing the scaling parameters appropriately, $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ will take different values so that $F^{\prime}(x)$ will be undefined on all bend points.
+Figure 10 highlights some important properties about composing triangle functions.
+Peaks alternate with valleys.
+Peak locations in one layer become valleys in the next.
+Valleys in one layer remain valleys in all future layers since 0 is a fixed point of each $T_{i}$.
+To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$.
+Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition.
+On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected.
+Due to the reflection, the slopes of $W_{i}$ on each side of a peak or valley are proportional.
+Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$.
+Either way, it’s important to note that the slopes on each side of a bend scale identically during each composition.
+Before we begin reasoning about $F^{\prime}(x)$, it can simplify the analysis to only consider the derivative of the error function $E^{\prime}(x)$.
+Lemma A.1.
+for $x\in P_{i}$, $F^{\prime}(x)$ is defined if and only if $E^{\prime}_{i}(x)$ is defined.
+Proof.
+All of $W_{n}(x)$ for $n<i$ are differentiable at $x\in P_{i}$ since $x$ will lie in the interior of a linear region of $W_{n}$.
+Therefore, $f^{\prime}_{i}(x)=\sum_{n=0}^{i-1}s_{n}W^{\prime}_{n}(x)$ exists at these points.
+Since $E^{\prime}_{i}+f^{\prime}_{i}=F^{\prime}$, $F^{\prime}(x)$ is defined if and only if $E^{\prime}_{i}(x)$ is defined.
+∎
+Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function at a point $x$. The left derivative will only be different by a constant factor.
+Lemma A.2.
+For all $x\in P_{i}$, $E_{+}^{\prime}(x)$ and $E_{-}^{\prime}(x)$ are proportional to
+$s_{i}-\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$ (5)
+Proof.
+Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in.
+To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$.
+The terms in the summation Equation 5 mostly derive from the chain rule.
+We will use $R$ to represent $W_{i+}^{\prime}(x)$.
+The first term in the sum will be $Rs_{i}$.
+$W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply.
+There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$.
+This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$.
+Note that the second term has the opposite sign as the first.
+For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$.
+For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$.
+Since this slope is positive, every term has the opposite sign as the first.
+Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula.
+Note that this same reasoning applies to the left side, the initial slope $L$ will just be different.
+∎
+Lemma A.3.
+If $E_{+}^{\prime}(x)=E_{-}^{\prime}(x)$, $E^{\prime}(x)$ must be equal to 0.
+Proof.
+Let $S$ represent Equation 5, and $R$ and $L$ be the constants of proportionality for the directional derivatives.
+If $E^{\prime}_{+}=E^{\prime}_{-}$, then $R_{x}S=L_{x}S$ for all $x\in P_{i}$.
+Since $W_{i}$ is comprised of alternating positive and negatively sloped line segments, $R_{x}$ and $L_{x}$ have opposite signs.
+The only way to satisfy the equation then is if $S=0$.
+Consequently, $E^{\prime}(x)=0$ for all $x\in P_{i}$.
+∎
+This lemma shows that to calculate the derivative at of $F(x)$ for any bend point $x$ (whenever it’s actually defined), one needs only to compute the derivative of the finite approximation $f_{i}$ (which excludes $W_{i}$).
+This will be useful later for proving other results.
+A visualization of this lemma is provided in Figure 11
+Lemma A.4.
+For all $x\in P_{i}$:
+$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (6)
+Proof.
+From the previous lemma we have $E^{\prime}(x)=0$ whenever the directional derivatives are equal.
+$F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$.
+The first $i-1$ terms are differentiable at the points $P_{i}$ since those points lie between the discontinuities in $B_{i-1}$.
+Therefore $F^{\prime}(x)$ is defined and can be calculated using the finite sum.
+∎
+We now prove our main theorem.
+Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions.
+The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks.
+Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.
+Theorem (3.1).
+$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
+$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$
+Proof.
+Rewriting Equation 5 (which is equal to 0) for layers $i$ and $i+1$ in the following way:
+$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1}{a_{i+2}}\left(s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}\right)$
+$s_{i+1}(1-a_{i+2})=s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}$
+allows for a substitution to eliminate the infinite sum
+$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}$
+Collecting all the terms gives
+$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}}$
+which simplifies to the desired result. ∎
+A.3 Sufficiency for differentiability
+We can show that in addition to being necessary for the existence of the derivative on the bend points, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
+Theorem A.5.
+$F^{\prime}(x)$ is well-defined if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$ and choose all $s_{i}$ according to Equation 4.
+Proof.
+We begin by considering Equation 5 for layer $i$ (it equals 0 by Theorem A.3).
+$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
+We will prove our result by substituting Equation 4 into this formula, and then verifying that the resulting equation is valid.
+First we would like to rewrite each $s_{n}$ in terms of $s_{i}$.
+Equation 4 gives a recurrence relation.
+Converting it to an non-recursive representation we have:
+$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (7)
+When we substitute this into Equation 5, three things happen: each term is divisible by $s_{i}$ so $s_{i}$ cancels out, every factor in the product except the last cancels, and $1-a_{i+1}$ cancels.
+This leaves
+$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (8)
+This equation has an interpretation that is important to the argument.
+1 is the full size of the initial discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added.
+Every time a wave is added, it can be thought of as subtracting from each side the first term appearing on the right.
+We will now show that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error.
+Inductively we can show:
+$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (9)
+This means that as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
+As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$.
+Assuming the above equation holds for all previous values of $j$
+$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
+using the inductive hypothesis to make the substitution
+$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
+Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded by the exponentially decaying series $(1-c)^{n}$, which approaches zero. ∎
+A.4 Error Decay
+Lemma A.6.
+The ratio $s_{i+2}/s_{i}$ is at most $0.25$.
+Proof.
+by applying Equation 4 twice, we have
+$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
+To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
+Since each $W_{i}$ takes values between 0 and 1, its contribution to $F$ is bounded by $s_{i}$.
+Since the $s_{i}$ decay exponentially, one could construct a geometric series to bound the error of the approximation and arrive at an exponential rate of decay.
+A.5 Second Derivatives
+Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a second derivative defined at the bend locations.
+To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the limits of these series on the right and left are not the same (unless all $a_{i}=0.5$), which implies that $F^{\prime\prime}(x)$ does not exist (see Figure 11 below).
+First we will produce the series of $\Delta x$.
+Let $x$ be the location of a peak of $W_{i}$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
+Lemma A.7.
+If $c<a_{i}<1-c$ for all $i$, we have $\lim_{n\rightarrow\infty}r_{n}=\lim_{n\rightarrow\infty}l_{n}=x$. Furthermore, $r_{n},l_{n}\neq x$ for any finite $i$.
+Proof.
+Let $R$ and $L$ denote the magnitude of $W^{\prime}_{i}$ on the left and right of $x$.
+$x$ is a peak location of $W_{i}$, so the right side slope is negative and the left is positive.
+Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})/L$ and $r_{1}=x+(1-a_{i+1})/R$.
+On each subsequent iteration $i+n$ ($n\geq 2$), $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$.
+Since $x$ is a valley point the right slope is positive and the left is negative.
+The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $W_{i+n}$ ranges from 0 to 1 over these spans.
+Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$.
+The resulting non-recursive formulas are:
+$x-l_{n}=\frac{1-a_{i+1}}{L}\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{1-a_{i+1}}{R}\prod_{m=2}^{n}a_{i+m}$ (10)
+The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
+Figure 11: The derivatives of the first few stages of approximation.
+Notice that each time a constant segment“splits”, the two neighboring segments adjacent to the split converge back to the original value.
+The points in the series approaching $x$ from the right are marked.
+$a_{0}\neq 0.5$ and all other parameters are set to $0.5$, which will cause the left and right sets of points to lie on lines with different slopes.
+Next we derive the values of $\Delta y$ to complete the proof.
+Theorem A.8.
+$F(x)$ cannot be twice differentiable unless $F(x)=x^{2}$.
+Proof.
+The points $l_{n}$ and $r_{n}$ are all peak locations, Equation 6 gives their derivative values as $f^{\prime}_{i+n}(r_{n})$.
+In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$.
+Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $R*s_{i}$ rather than 1.
+Focusing on the right hand side we get:
+$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=R*s_{i}\prod_{m=2}^{n}(1-a_{i+m})$
+taking $\Delta y/\Delta x$ gives a series:
+$\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
+The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$.
+The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$.
+Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$.
+∎
+A.6 Concavity
+Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure11).
+This then allows us to show that $F$ is concave.
+Lemma A.9.
+The function $F(x)$ is concave when all $s_{i}$ are chosen according to Equation 4.
+Proof.
+To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i+1}[2^{i}]]$ for $0\leq n\leq 2^{i}$, which tracks the values of $F^{\prime}$ at the $i$th set of valley points.
+All but the first and last points will have been peaks at some point in their history, so Equation 6 gives the value of those derivatives as $f^{\prime}_{i}$.
+We establish two inductive invariants.
+One is that the y-values in the list $Y_{i}$ remain sorted in descending order.
+The other is that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$, indicating that the constant value segments of $f_{i}$ lie in between the limits in the list $Y_{i}$.
+Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave.
+Which then can prove that their limit $F$ is also concave.
+As the base case $f_{0}$ is a line with derivative 0, $V_{0}$ contains its two endpoints.
+$Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore both the points in $Y^{\prime}$ are in descending sorted order.
+The second part of the invariant is true since 0 is in between those values.
+Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value, which we know from the second inductive hypothesis is in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$.
+The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$.
+This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order of $Y^{\prime}$.
+Adding $s_{i}W_{i}$ takes $f_{i}$ to $f_{i}+1$ splitting each constant valued interval in two about the points $P_{i}$, increasing the left side, and decreasing the right side.
+Recalling from the derivation of Equation 5 all terms but the first in the sum have the same sign, so the limiting values in $Y^{\prime}_{i}$ are approached monotonically.
+Using the first inductive hypothesis, we have on the left interval $Y^{\prime}_{i}[n]=Y^{\prime}_{i+1}[2n]>f^{\prime}_{i+1}>f^{\prime}_{i}=Y^{\prime}_{i+1}[2n+1]$ and on the right we have $f^{\prime}_{i}=Y^{\prime}_{i+1}[2n+1]>f_{i+1}>Y^{\prime}_{i}[n+1]=Y^{\prime}_{i}[2n+2]$.
+And so each constant interval $f_{i+1}$ remains bounded by the limits in $Y^{\prime}_{i+1}$.
+We will now show by contradiction that the limit $F$ of the sequence of concave $f_{i}$ is also concave.
+Assume that $F$ is non-concave.
+Then there exist points $a$, $b$, and $c$ such that $F(b)$ lies strictly below the line connecting the points $(a,F(a))$ and $(c,F(c))$.
+Lets say it’s below the line by an amount $\epsilon$.
+Since at each point $f_{i}$ converges to $F$, we can find $i_{a}$ such that $f_{ia}(a)-F(a)<\epsilon/2$, etc…, we take $i=\max(i_{a},i_{b},i_{c})$.
+Since $f_{i}(a)$ and $f_{i}(c)$ are no more than $\epsilon/2$ lower than their limiting values, the entire line connecting $(a,f_{i}(a))$ and $(c,f_{i}(c))$ is no more than $\epsilon/2$ lower than the line between $(a,F(a))$ and $(c,F(c))$.
+$f_{i}(b)$ is also no more than $\epsilon/2$ higher than $F(b)$, thus $f_{i}(b)$ must still lie below the line between $(a,f_{i}(a))$ and $(c,f_{i}(c))$, making $f_{i}$ non-concave and producing a contradiction.
+∎
+A.7 Learning Rates
+Figure 12: Learning $y=x^{3}$ at various learning rates for 1000 epochs.
+The figure shows the log losses of the best model of 30 for networks trained according to the methods in this paper versus random initialization and gradient descent (“Default Network”).
+For high learning rates, neither learns because the steps are too large.
+For low learning rates (below $10^{-5}$), the steps are too small, and our networks are likely deriving an implicit advantage from being forcibly initialized to a convex function.
+Both methods are able to converge for learning rates in between $10^{-4}$ and $10^{-2}$; one could run for more epochs to see a similar advantage of our method for smaller learning rates, if desired.
+All results in the main body of the paper used a constant learning rate of $10^{-3}$.
+In this appendix, we considered an ablation study on the learning rate for the task of learning $y=x^{3}$.
+As seen in Figure 12, the learning rate we selected was approximately optimal for both our method as well as default network training.
+We note that for this ablation study, a constant 1000 epochs were run, which explains why both methods perform worse as the learning rate becomes minuscule.
+At small learning rates, what is really measured is how close of a guess the initialization is to the target function.
+Our networks are preforming better here simply because they are always outputting convex functions.
+But this only accounts for losses on the order of $10^{-4}$, which indicates that in the more reasonable learning rate ranges, our pretraining is performing a meaningful function and enabling order-of-magnitude improvements over default network training.
+Potential Broader Impact
+This paper presents work whose goal is to enable more efficient neural networks.
+While the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.

### v5

@@
 However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space.
-The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.
+The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] See Appendix A.2 for definitions of terms like linear regions, activation patterns, and activation regions.
 and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018).
@@
 Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size.
-Although the preliminary results in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
+Although preliminary numerical demonstrations in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
 2 Related Work
 This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component.
-In one dimension, where we initially focus, our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section.
-The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
+In one dimension, where we initially focus, our work is a generalization of the approximation to $x^{2}$ we review in this section.
+The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
+The goal of this method is not to re-express the full set of neural network weights, but to constrain them into efficient patterns.
 Figure 1: (Top to bottom) Composed triangle waves; using collections of the above function to approximate $x^{2}$; derivatives of the above approximations.
-2.1 Related Work in Approximation
+2.1 Function Approximation
 Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989).
 Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019).
-In finite cases, the trade-off between width and depth is often studied in assessing a network’s ability to approximate (learn) a function.
+In finite cases, one may study trade-offs between width and depth to assess a network’s ability to approximate (learn) a function.
 Notably, there exist functions that can be represented with a sub-exponential number of neurons in a deep architecture, yet which require an exponential number of neurons in a wide and shallow architecture.
@@
 But these problems can be rectified by taking a sum over the layers of a network.
-Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves.
+Yarotsky (2017) and Liang & Srikant (2016) construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves.
 To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1.
@@
 The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems.
-One possible use case is to construct a multiplication gate.
 Perekrestenko et al.
-(2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$.
+(2018) construct a multiplication gate via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$.
 The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms.
@@
 In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
-Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022).
+Other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022), including spline-type methods (Eckle & Schmidt-Hieber, 2019).
 Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present.
-Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks.
-It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
-Finally, Park et al.
+Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize sigmoid neural networks for classification.
+Similar to the spirit of our work, which restricts parameterizations of ReLU networks, Elbrächter et al.
+(2019) explores theoretical aspects of the conditionining of ReLU network training and provides constructive results for a parameterization space that is well-conditioned.
+Chen & Ge (2024) present a creative approach where they explore reparameterizing the direction of weight vectors using hyperspherical coordinates to improve training dynamics.
+Unlike their reparameterization, ours will restrict the network’s expressivity in order to prevent it from learning inefficient weight patterns.
+Lastly, Park et al.
 (2021) approaches the problem of linear region maximization from an information theory perspective and uses a loss penalty rather than a reparameterization to increase the number of linear regions.
-2.2 Related Work in Initialization
+2.2 Neural Network Initialization
 Our work seeks to improve network initialization by making use of explicit theoretical constructs.
@@
 They use weight values that are sampled from distributions defined by the input and output dimension of each layer.
-Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization.
+Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value.
+The ReLU activation then sets the output of that layer to 0, blocking any gradient updates.
 This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023).
-It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value.
-The ReLU activation then sets the output of that layer to 0, blocking any gradient updates.
 Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020).
@@
 Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion.
-In this way, we can avoid having to use the extra intermediate neurons.
+In this way, we can avoid having to use extra intermediate neurons.
 This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
@@
 3.1 Differentiable Model Output
-Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged or fractal with many choices of scaling parameters.
-This can lead to the network settling into solutions that are not predictive of the areas in between training points (or possibly not even interpolating the training data) due to rapid oscillation.
+Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters.
+This would be poorly predictive of unseen data points generated from a smooth curve.
 Our main mathematical result addresses this issue by forcing differentiability of the network output.
-To achieve differentiability, it turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them:
+We do so by making the sum of the derivatives of the triangle waves converge.
+For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
+For a broader mathematical discussion, as well as sufficient conditions for differentiability in the limit, see the appendix.
 Theorem 3.1.
-$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
+$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$ and is continuous on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
 $s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (4)
@@
 Accordingly, in this section, we implement our method to learn several convex one-dimensional curves.
-These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6.
+These nonlinear functions are known—so that we can measure error—and we remark and demonstrate results on more difficult functions in Section 5, Section 6, and Appendix A.10.
 The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
@@
 We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability.
-The intent of both comparisons is see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape.
+The intent of both comparisons is to see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape.
 Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@
 Training Type Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
-Default Network $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
+Default Network (Kaiming) $2.11\times 10^{-5}$ $2.19\times 10^{-5}$ $7.20\times 10^{-2}$ $2.82\times 10^{-2}$
 RAAI Distribution $2.14\times 10^{-5}$ $4.40\times 10^{-5}$ $3.97\times 10^{-2}$ $4.12\times 10^{-2}$
@@
 Training Type Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
-Default Network $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
+Default Network (Kaiming) $4.50\times 10^{-5}$ $5.75\times 10^{-5}$ $1.15\times 10^{-1}$ $1.96\times 10^{-1}$
 RAAI Distribution $3.59\times 10^{-5}$ $1.09\times 10^{-5}$ $3.63\times 10^{-2}$ $2.31\times 10^{-2}$
@@
 Training Type Min $x^{3}$ Min $x^{11}$ Min $\sin(x)$ Min $\tanh(3x)$
-Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
+Default Network (Kaiming) $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
 Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
 Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends.
-First, the worst performing networks are those that rely on purely randomized initializations.
+First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization.
 Even the networks that forgo pretraining benefit from initializing with many activation regions.
@@
 RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
-Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
+Figure 4: Comparison between standard Kaiming initialization/gradient descent and pretraining with differentiablity enforced.
+Using more linear regions allows the curve to better predict the test points.
 Our second set of results is shown in Table 3 and in Figure 4.
@@
 (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions).
-We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising (and of course, assembling more copies of our network into a deeper or wider larger network would give additional representational power).
+We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising.
+In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world datasets and problems with up to eight input dimensions.
 We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1.
@@
 Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks.
-Rather than directly using our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect).
+Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect).
 Two slight modifications are made to each of the component networks to better interact with ReLU.
@@
 We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture.
-We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU or even other networks.
+We note that there are other metrics for evaluating a neural network’s expressivity (e.g., Raghu et al.
+(2017)), which would be interesting to measure for our networks in future work.
+Overall, we are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU and other classes of neural networks.
 Appendix A Appendix
-Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks, as creating nonconvex functions this way will tend to make a lot of negative values.
+Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks. Colors correspond to Figure 2.
 A.1 Initialization and Training Algorithm
@@
 end while
-The network weights can then be initialized once more based on the learned vector $A$ so that the network weights can be updated via regular gradient descent.
+The network weights can then be initialized once more based on the learned vector $A$, to then update the weights via regular gradient descent.
 In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs.
 Full pseudocode is listed in Algorithm 1.
-A.2 Derivations
-For convenience we first restate the functions defined in the main body of the paper.
+A.2 Terminology
+Before presenting further formal mathematical details of our method, we first briefly review a few pieces of basic ReLU network terminology used in the paper.
+The reader is referred to the Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
+An activation pattern is a boolean mask that tracks which neurons in a network have their output zeroed by ReLU activations.
+The activation regions of a ReLU network are connected (and can be shown to be convex) sets of inputs on which the activation pattern is constant.
+Since the action of the ReLU activation function is constant, the network output over an activation region is equivalent to the case where there is no activation function and the associated zeroed neurons are absent; therefore, the network output behaves linearly over the inputs in the activation region.
+Relatedly, a linear region is a set of inputs on which the network output behaves linearly with respect to its inputs.
+It may consist of multiple neighboring activation regions.
+Another important concept is the boundary of a neuron, as described in Rolnick & Kording (2020): the set of inputs for which the neuron outputs 0, independently of ReLU.
+The boundary of a neuron is precisely the boundary of the activation regions it adds to the network.
+Chen & Ge (2024) refers to this set as the characteristic activation boundary since these are the boundaries of the activation regions.
+We note that although works like ours are theoretical papers that leverage these concepts, studying these ideas can lead to interesting applied learning research.
+For instance, Rolnick & Kording (2020) leverages theoretical works like those cited in our related work discussion for an exciting privacy application: it is proven that a ReLU network’s output can often be provably used to reverse engineer the architecture of a ReLU network, up to isomorphism.
+A.3 Mathematical Results
+For convenience, we first restate the functions defined in the main body of the paper.
 $T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases}$
@@
 $F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$
+The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$.
+We begin by assuming that
+$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$
+We will see that the resulting choice of $s_{i}$ ensures uniform convergence of the derivative terms, so that the derivative of the infinite sum is indeed infinite sum of the derivatives.
+Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right respectively.
+The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
 Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$.
@@
 The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network.
-Although the derivative of an infinite sum of functions is not always equal to the sum of the individual functions derivatives, it is in this case, so notationally we use $F^{\prime}$ to describe the latter for brevity.
+This odd split around layer $i$ makes the proofs cleaner.
 Figure 10: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
-The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and on all of $[0,1]$.
-Fortunately, the left and right derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist, since each bend in each $W_{i}$ will use the slope of the line segment to its left or right respectively.
-The problem is that without choosing the scaling parameters appropriately, $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ will take different values so that $F^{\prime}(x)$ will be undefined on all bend points.
 Figure 10 highlights some important properties about composing triangle functions.
@@
 ∎
-Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function at a point $x$. The left derivative will only be different by a constant factor.
+Thanks to the previous lemma, we only need to work with $E_{i}$.
+Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function at a point $x$.
+The left derivative will only be different by a constant factor.
 Lemma A.2.
@@
 Proof.
-Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in.
-To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$.
-The terms in the summation Equation 5 mostly derive from the chain rule.
-We will use $R$ to represent $W_{i+}^{\prime}(x)$.
-The first term in the sum will be $Rs_{i}$.
-$W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply.
-There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$.
-This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$.
+Let $x_{k}$ be some point in $P_{i}$, and let $k$ be its index in any list it appears in.
+To calculate the value of $E_{+}^{\prime}(x_{k})=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x_{k})$, we will have to find the slope of the linear intervals to the immediate right of $x_{k}$ for all $W_{i}$.
+We will use $R_{x}$ to represent $W_{i+}^{\prime}(x_{k})$.
+The first term in the sum will be $R_{x}s_{i}$.
+Since the derivatives of composed functions will multiply from the chain rule, so the value of the next term is $W^{\prime}_{i+1}(x_{k})=T^{\prime}_{i+1}(W_{i}(x_{k}))R_{x}$.
+$T_{i+1}$ has two linear segments, giving two slope possibilities to multiply by.
+The correct one to choose is $-1/(1-a_{i+1})$ because it’s ‘active’ around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$).
+This gives $W^{\prime}_{i+1}(x_{k})=-R_{x}\frac{s_{i+1}}{(1-a_{i+1})}$.
 Note that the second term has the opposite sign as the first.
-For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$.
-For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$.
-Since this slope is positive, every term has the opposite sign as the first.
-Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula.
-Note that this same reasoning applies to the left side, the initial slope $L$ will just be different.
+For all remaining terms, since $x_{k}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$.
+For $x\in(B_{j+1}[k-1],B_{j+1}[k+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the slope $1/a_{j+1}$.
+Since this slope is positive, every remaining term continues to have the opposite sign as the first term.
+Summing up all the terms with the coefficients $s_{i}$, and factoring out $R_{x}$ will yield the desired formula.
+Note that this same derivation applies to the left sided derivatives as well because the ‘active’ slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right.
+The initial slope constant $L_{x}$ will just be different.
 ∎
@@
 ∎
-This lemma shows that to calculate the derivative at of $F(x)$ for any bend point $x$ (whenever it’s actually defined), one needs only to compute the derivative of the finite approximation $f_{i}$ (which excludes $W_{i}$).
+The following lemma shows that to calculate the derivative at of $F(x)$ for any bend point $x$, one needs only to compute the derivative of the finite approximation $f_{i}$ (which excludes $W_{i}$).
 This will be useful later for proving other results.
-A visualization of this lemma is provided in Figure 11
 Lemma A.4.
@@
 Therefore $F^{\prime}(x)$ is defined and can be calculated using the finite sum.
+A visualization of this lemma is provided in Figure 11.
 ∎
-We now prove our main theorem.
-Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions.
+We now prove our main theorem, which shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting approximation converges to a continuously differentiable function.
 The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks.
-Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.
+Once they are both valleys, they will be treated the same by the remaining compositions, so the sizes of their remaining discontinuities will need to be proportional.
 Theorem (3.1).
-$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
+$F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
 $s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$
@@
 which simplifies to the desired result. ∎
-A.3 Sufficiency for differentiability
-We can show that in addition to being necessary for the existence of the derivative on the bend points, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
+Figure 11: The derivatives of the first few stages of approximation.
+Notice that each time a constant segment “splits”, the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it’s from)
+A.4 Sufficiency for differentiability
+This form of the scaling equation derived in the previous section is rather interesting.
+Since the ratio of two successive scaling terms is $(1-a_{i+1})a_{i+2}$, factors of both $a_{i}$ and $1-a_{i}$ are present in $s_{i}$.
+This has a few implications, firstly if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth.
+Secondly, having both $a_{i}$ and $1-a_{i}$ will cancel whichever slope multiplier $T_{i}$ contributes to $W_{i}$ at each point $x$, leaving behind the other term, which is less than 1.
+If we ensure each $a_{i}$ is bounded away from 0 or 1 by being drawn from the an interval such that $c<a_{i}<1-c$ for some $0.50<c<$ 0.5, then the maximum value of $W^{\prime}_{i}$ is always upper bounded by $(1-c)^{i}$, which is sufficient for uniform convergence.
+This condition also means that the bends points (and thus the activation regions of the corresponding neural network) will become dense in $[0,1]$, as each region is partitioned (at worst) in a $c$ : $1-c$ ratio.
+Lastly, we can show that bounds on $a_{i}$ and our choice of scaling values are sufficient for the existence of $F^{\prime}$ on the bend points, in addition to being necessary for the existence of the derivative on the bend points, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
 Theorem A.5.
-$F^{\prime}(x)$ is well-defined if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$ and choose all $s_{i}$ according to Equation 4.
+on bend points $x$, $F^{\prime}(x)$ exists if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$ and choose all $s_{i}$ according to Equation 4.
 Proof.
@@
 We will prove our result by substituting Equation 4 into this formula, and then verifying that the resulting equation is valid.
-First we would like to rewrite each $s_{n}$ in terms of $s_{i}$.
+First we would like to rewrite each occurrence of $s$ in terms of $s_{i}$.
 Equation 4 gives a recurrence relation.
@@
 $1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (8)
-This equation has an interpretation that is important to the argument.
-1 is the full size of the initial discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added.
-Every time a wave is added, it can be thought of as subtracting from each side the first term appearing on the right.
-We will now show that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error.
+This equation has a meaningful interpretation that is important to the argument.
+1 is the full size of the initial derivative discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added.
+Every time a wave is added, it subtracts the first term appearing on the right hand side.
+The following argument shows that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining discontinuity, guaranteeing its disappearance in the limit.
 Inductively we can show:
 $1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (9)
-This means that as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
+In words this means that as the first term appearing on the right in Equation 8 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
 As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$.
@@
 Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded by the exponentially decaying series $(1-c)^{n}$, which approaches zero. ∎
-A.4 Error Decay
+A.5 Error Decay
 Lemma A.6.
@@
 Since the $s_{i}$ decay exponentially, one could construct a geometric series to bound the error of the approximation and arrive at an exponential rate of decay.
-A.5 Second Derivatives
-Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a second derivative defined at the bend locations.
+A.6 Second Derivatives
+Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a continuous second derivative, as it will not be defined at the bend locations.
 To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the limits of these series on the right and left are not the same (unless all $a_{i}=0.5$), which implies that $F^{\prime\prime}(x)$ does not exist (see Figure 12 below).
@@
 The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
-Figure 11: The derivatives of the first few stages of approximation.
-Notice that each time a constant segment“splits”, the two neighboring segments adjacent to the split converge back to the original value.
+Figure 12: An illustration of attempting to calculate the second derivative.
 The points in the series approaching $x$ from the right are marked.
+We rely on the fact that at bend points, the first derivative converges back to the value it had at a finite point in the approximation.
 $a_{0}\neq 0.5$ and all other parameters are set to $0.5$, which will cause the left and right sets of points to lie on lines with different slopes.
@@
 In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$.
-Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $R*s_{i}$ rather than 1.
+Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $Rs_{i}$ rather than 1.
 Focusing on the right hand side we get:
@@
 $\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
-The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$.
-The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$.
-Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$.
+The issue which arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$.
+The only way for these formulas to agree then is for $R^{2}=L^{2}$ which implies $a_{i}=1-a_{i}=0.5$.
+Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$).
 ∎
-A.6 Concavity
+A.7 Monotonicity and Continuity of Derivatives
 Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure11).
-This then allows us to show that $F$ is concave.
+This can extend into the limit to show that $F^{\prime}$ is monotone decreasing and that $F$ is concave.
 Lemma A.9.
@@
 Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave.
-Which then can prove that their limit $F$ is also concave.
+Which then be used to prove that their limit $F$ is also concave.
 As the base case $f_{0}$ is a line with derivative 0, $V_{0}$ contains its two endpoints.
@@
 ∎
-A.7 Learning Rates
+Lastly we briefly sketch out why $F^{\prime}$ is continuous.
+It relies on some of our earlier reasoning.
+Monotonicity of the derivative makes continuity easy to show, because when $x_{1}$ within $\delta$ of $x_{2}$ has $f(x_{1})$ within $\epsilon$ of $f(x_{2})$, so do all intermediate values of $x$.
+We can establish continuity of $F^{\prime}$ at bend points $x$ easily by using Equation 9.
+We can pick an $i$ so that the constant value segments of $f^{\prime}_{i}$ are within $\epsilon$ of $F^{\prime}(x)$ and then use the next iteration of bend points (since the constant intervals split, but the new segments near $x_{n}$ converge monotonically towards it) to find $\delta$.
+In the case of continuity for non bend points $x$, they sit inside a constant-valued interval of $f_{i}$ for each $i$.
+We can choose $i$ such that $\sum_{n=i}^{\infty}(1-c)^{n}<\epsilon/2$ because this series constrains how far derivative values can move in the limit, and then use the constant interval $x$ is situated in to find $\delta$.
+A.8 Activation Region Counting and Zaslavsky’s Theorem
+In the coming sections, we present the results of using our 4-neuron-wide construction as an activation function on more complicated datasets.
+It will be useful to know the impact of this approach on the activation region count when assessing how to best spend a network’s parameter budget.
+To build up to it, we will provide an informal introduction of Zaslavsky’s theorem (Zaslavsky, 1975).
+The activation boundaries of ReLU neurons in a single hidden layer network form a hyperplane arrangement.
+Counting the number of cells, which are convex connected components of space separated by the hyperplanes, is equivalent to determining the number of activation regions produced by the shallow ReLU network.
+A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc… General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network).
+Zaslavsky’s theorem states that
+Theorem A.10.
+When in general configuration, the number of cells ($R\binom{n}{D}$) created by a $D$-dimensional arrangement of $n$ hyperplanes is
+$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}$
+To compute this sum, one would go down to the row of Pascal’s triangle corresponding to the number of planes $n$, and sum the first $D$ entries (see Figure 13 for a visual example).
+It’s a known fact that the sequence of the $d$th entry for each row grows to the $d$th power.
+So as the number of hyperplanes tends towards infinity, this sum is dominated by the $D$th term, which grows to the power $D$.
+In other words, the number of linear regions formed by a single hidden layer ReLU network grows exponentially with respect to its input dimension.
+This property partly explains why neural networks are successful at learning high dimensional functions, even when other classical methods become intractable.
+Figure 13: 3 lines dividing 2-dimensional space creates 7 cells.
+This is calculated by going down to row 3 (for the number of lines) and summing the numbers at indices 0, 1, and 2 (for the dimension).
+Also note that there is one 2-D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
+Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e.
+points, lines, planes, etc…) are given by the entries in Pascal’s triangle.
+So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower dimensional objects, or that each lower dimensional object could be ‘assigned’ a cell.
+There’s a striking similarity in the intuitions behind the recurrence that generates pascal’s triangle, and the calculation of the numbers of lower dimensional objects created from a hyperplane arrangement.
+In the case of Pascal’s triangle, if one wished to know how to choose 3 items from 6, one could reason that if item 6 is included, 2 items would be chosen from the remaining 5, and if item 6 is excluded, 3 would be chosen from the remaining 5, and therefore $\binom{6}{3}=\binom{5}{2}+\binom{5}{3}$.
+Now consider the case of 3 planes partitioning 3-space, and the addition of a 4th plane.
+The total number of lower dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points).
+One could also picture that the new 4th plane ‘cuts’ the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 13), and count the 2-cells that appear.
+Either reasoning leads to $R\binom{n}{D}=R\binom{n-1}{D}+R\binom{n-1}{D-1}$.
+This similarity between the recurrences means that if the initial conditions can be lined up suitably, Zaslavsky’s theorem can be proven by induction over Pascal’s triangle.
+In the case of our architecture, using our 1-dimensional construction as an activation function will generate $n$ generally configured sets of $m$ parallel hyperplanes.
+To account for this, a slight modification is made to the recurrence.
+Adding an additional set of $m$ hyperplanes would add $m$ times as many regions, as each of the lower dimensional objects in the existing configuration would project onto each of the $m$ additional hyperplanes.
+The effect that this has on the formula is that $R\binom{n,m}{D}=R\binom{n-1,m}{D}+m\times R\binom{n-1,m}{D-1}$.
+Inducting over Pascal’s triangle (with a base case of $R\binom{n,m}{1}=1+nm$) would give the non-recursive formula of $R\binom{n,m}{D}=\sum_{d=0}^{D}m^{d}\binom{n}{d}$.
+This means that out architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic trade off once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations).
+Of course, the effect on accuracy will all depend on how well these regions can be utilized.
+A.9 Learning Rates
 Figure 14: Learning $y=x^{3}$ at various learning rates for 1000 epochs.
@@
 But this only accounts for losses on the order of $10^{-4}$, which indicates that in the more reasonable learning rate ranges, our pretraining is performing a meaningful function and enabling order-of-magnitude improvements over default network training.
+A.10 Real-World Data and Classification Problems
+Here we present a few preliminary results on extending our pretraining technique to classification problems and real-world datasets.
+The classification problem we chose is the classic two spirals dataset, and we selected the UCI dataset “Concrete Compressive Strength” (Yeh, 1998) for our real-world regression task.
+The concrete dataset has 8 numerical features that can be used to predict the compressive strength of a concrete sample.
+The networks are set up as described before, with our 4-neuron-wide networks acting as 1D to 1D activation functions inside randomly initialized standard linear layers.
+In these experiments, we use one “hidden layer” of our networks (which we choose to have depth 5 in our tests).
+The geometric interpretation of this is that we are setting up one-dimensional convex functions oriented in random directions on the input space, and then taking a linear combination of them as the network output.
+In the case of classification, the loss function is simply swapped for cross entropy.
+After our pretraining phase, there are two choices of how to conduct the second phase of training (training matrix entries directly).
+All the parameters could be freed from constraints (“dense”), or the smaller 4-neuron subnetworks could be kept isolated from each other (but otherwise have their parameters freed).
+In the case where the 4-neuron networks remain in isolation, the weight matrices of the hidden layers will have a block diagonal structure (block size 4—the width of each subnetwork).
+Thus, we consider two fair (same number of free parameters) comparisons in this subsection: (1) dense versions of our and Kaiming-initialized networks, and (2) block diagonal versions of our and Kaiming-initialized networks.
+As discussed below and in the figure captions, we find that while the dense variant of our method ties or is slightly worse than a regular fully-connected network of the same dimensions, in the block diagonal case, our experimental networks significantly outperform their Kaiming-initialized counterparts.
+This makes sense in light of the experiments from the main body, where we can effectively shape the output of 4-neuron-wide networks better than random initializations, but where ensembles of our networks (see approximation of $x^{3}-x$ in Figure 7) get filled in with noise by gradient descent during the second training stage.
+This again highlights the need for more mathematical developments to provide a better extension into higher-dimensional nonconvex functions.
+Figure 15: Train/test loss plots for Kaiming-initialized (“ordinary”) fully-connected networks and our method (“experimental”).
+Our method has two steps: pretriaining on triangle peaks, followed by the optimization of the raw matrix weights.
+The switch-off is at epoch 500, hence the associated visible change in the loss curves.
+The block diagonal variants of the networks (right) are generally worse at the task (test losses 30.1 (ours) and 40.8 (Kaiming)) than the dense variants (losses of 22.3 (ours) and 21.2 (Kaiming)).
+Our experimental networks outperform by up to an order of magnitude in the block diagonal case.
+We used 32 of our convex blocks (i.e., all weight matrices are size 128, including for standard comparison networks).
+Figure 16: Class predictions of block diagonal networks on a standard two-spiral dataset; ours (left) and Kaiming-initialized (right).
+While the cross-entropy losses are comparable for the dense-matrix variants of both networks, when a block diagonal structure is imposed, Kaiming initialization fails to learn the spiral.
+Note that the colors of the network class predictions are inverted for visibility.
+Cross-entropy losses are 0.0032 and 0.63 respectively.
+We used 16 of our convex blocks, so all weight matrices are 64 by 64.
+Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 15 and 16 already show impressive results using our current approach.
+In the case of the concrete problem (Figure 15), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced.
+In the case of the two-spiral classification task (Figure 16), our network is able to learn an accurate decision boundary (left subfigure), whereas a standard network constrained to be block diagonal fails to learn (right subfigure).
+When dense weight matrices are trained, our networks are almost able to tie with the standard initializations, although convergence can take longer.
+These results suggest that our method might be quite powerful when its parameter budget is spent on greater width and depth of convex function blocks (as in Figure 9) instead of filling in dense matrices.
+Our two-step training approach adds some practical challenges, such as deciding when to switch parameterizations for optimal convergence time.
+Additionally, since the optimization variables are different, the optimizer will have to restart any momentum or adaptive learning rates when the switch is made, which can sometimes cause loss to temporarily spike.
+We lowered the learning rate on the second step to avoid this.
+Further interesting optimizations of our approach (e.g., training another network to inform our algorithm when to switch parameterizations) are imagined as future work.
+A.11 Large-Scale Image Classification Examples
+We include very preliminary demonstrations of how our method can increase performance on large-scale image classification tasks.
+A nice feature of our method is that it can be used anywhere dense layers appear in networks, which enables its use in CNN architectures.
+Nonetheless, combining our construction with convolutions or other layers is a use case far outside of the scope of some of the toy problems where we observe orders of magnitude of improvement, and the output of each subnetwork might not learn to be as nicely behaved as in the tightly controlled experimental setting of 1D convex training data.
+Thus, we stress that while our initial results are positive, and we can reasonably conjecture why, future work on extending our method to arbitrary ReLU networks would provide greater rigor and likely far greater performance gains, as seen with our lower-dimensional function approximation examples that realize the full benefit of our mathematically principled approach.
+We evaluated two standard benchmarks: CIFAR-10 and ImageNet.
+In the case of CIFAR-10, we trained a small network of 3 convolutions followed by 2 fully connected layers.
+To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added.
+In the case of ImageNet, we used PyTorch’s pretrained VGG16 model for its convolution weights (which were frozen during training).
+We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization.
+We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImagnNet.
+Figure 17: results for CIFAR-10 experiments.
+Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts; for instance, at epochs 3 and 4, which per the loss curves’ characteristics are where one would likely stop training.
+While the loss curves are similar in shape, we report that at epoch 4, for example, our training and validation losses are 25.6% and 5.1% lower than their standard network counterparts.
+Figure 18: In our ImageNet test we similarly see modest gains of our (“experimental”) network over a Kaiming-initialized (“ordinary”) network, yet due to using a smaller learning rate here ($0.00001$, vs.0.0010.0010.0010.001 as used with CIFAR-10), the loss curves are smoother, so we see our networks consistently outperforming the standard network 25% into epoch 1.
+We trained both networks for 2 epochs; at epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).
+These two examples show that even for large-scale image classification tasks—far away from the analytic 1D function tests considered in the main body—our methodology still has the potential to yield noticeable benefits, highlighting the promise of extending the present work to more general network architectures as a future research direction.
 Potential Broader Impact
 This paper presents work whose goal is to enable more efficient neural networks.
-While the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
+Although the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.

### v6

@@
 Abstract
-A neural network with ReLU activations may be viewed as a composition of piecewise linear functions.
-For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly.
-Therefore, randomly initialized models are often unnecessarily large, even when approximating simple functions.
-To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to exhibit a number of linear regions exponential in depth.
-Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights.
+In a neural network with ReLU activations, the number of piecewise linear regions in the output can grow exponentially with depth.
+However, this is highly unlikely to happen when the initial parameters are sampled randomly, which therefore often leads to the use of networks that are unnecessarily large.
+To address this problem, we introduce a novel parameterization of the network that restricts its weights so that a depth $d$ network produces exactly $2^{d}$ linear regions at initialization and maintains those regions throughout training under the parameterization.
 This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
-We further demonstrate how to extend our approach to multidimensional and non-convex functions, with similar benefits observed.
+We further demonstrate how to extend our approach to multidimensional and non-convex functions, allowing it to replace the dense layers in other networks; preliminary improvements are shown for image classification on CIFAR-10 and ImageNet.
+Machine Learning, ICML
 1 Introduction
@@
 Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions.
-Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized.
-They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training.
-As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize.
+Indeed, Hanin & Rolnick (2019) show a rather disappointing bound for randomly initialized networks: the average number of linear regions formed at initialization is invariant with depth.
+Whether using a deep or shallow architecture, only the total number of neurons matters.
+They further observed that gradient descent has a difficult time creating new activation regions so that their bounds approximately held after training.
+As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see 5).
 Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
-The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting.
-Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain.
-We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights.
-This allows the network to discover solutions that are more accurate and unlikely to be found otherwise.
-Our technique directly rectifies the issues raised by Hanin & Rolnick (2019).
-Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise.
-Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage.
-Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions.
-Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size.
-Although preliminary numerical demonstrations in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
+The present work presents mathematical algorithms that avoid the limitations of random initialization. Our contributions include:
+  • A reparameterization of a 4-neuron-wide, depth $d$ ReLU network that constrains it to initialize with $2^{d}$ activation regions over the input domain (Section 3)
+  • A novel pretraining strategy, which enforces the existence of $2^{d}$ activation regions during optimization to produce its initial solution (Sections 3, 3.1)
+  • Numerical results for one-dimensional test cases that yield orders-of-magnitude improvements in network performance (Section 4)
+  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Section 5; Appendices A.10A.11)
+While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and quickly outpacing such networks’ performance on more practical test cases such as ImageNet (Appendix A.11).
+We emphasize that while the preliminary mathematical analysis in this work pertains to a specific ReLU construction, this construction can be repeated and combined in useful ways to obtain performance improvements on larger networks (Section 5).
+We conclude in Section 6 with discussions on extending the mathematical framework of our method to arbitrary ReLU networks and other types of networks, which would have significant practical utility.
 2 Related Work
-This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component.
-In one dimension, where we initially focus, our work is a generalization of the approximation to $x^{2}$ we review in this section.
-The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
-The goal of this method is not to re-express the full set of neural network weights, but to constrain them into efficient patterns.
+This work is primarily concerned with a novel training methodology for ReLU networks, but its development stems from a more abstract idea of an exponentially accurate approximation to the function $x^{2}$, which can be encoded into a ReLU network.
+Our work generalizes this construction into a trainable family of differentiable convex one-dimensional functions (see Appendix A.3).
 Figure 1: (Top to bottom) Composed triangle waves; using collections of the above function to approximate $x^{2}$; derivatives of the above approximations.
@@
 The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)).
-Our reparameterization strategy focuses on non-symmetric triangle waves.
-The location on $(0,1)$ of a triangular function’s peak acts as an alternative parameterization of the function, instead of its weight matrix representation.
-Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions.
-Furthermore, this parameterization is trainable.
-By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions.
+Our reparameterization strategy (Section 3) focuses on non-symmetric triangle waves.
 The dilated triangular waveforms produced in this manner are not particularly useful on their own.
-Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit).
+Their oscillations quickly become excessively rapid, and their derivatives do not exist everywhere (especially in the infinite-depth limit).
 But these problems can be rectified by taking a sum over the layers of a network.
@@
 Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010).
-They use weight values that are sampled from distributions defined by the input and output dimension of each layer.
+They use weight values that are randomly sampled from distributions defined by the input and output dimension of each layer.
 Aside from sub-optimal approximation power associated with random weights, a common issue is that the initial weights and biases in a ReLU network can cause every neuron in a particular layer to output a negative value.
@@
 Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed.
-We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU.
+We observed during our experiments that RAAI greatly reduces, but does not eliminate, the likelihood of dying ReLU.
 Our approach enforces a specific network structure that does not collapse in this manner.
 3 Initialization and Pretraining Construction
-We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$.
-As alluded to in Section 2.1, our initialization algorithm is simply to associate a nonsymmetric triangle function with each layer, and then to randomize the location of each triangle’s peak in $(0,1)$; any choice produces $2^{d}$ regions.
-This yields a reparameterization of the ReLU network in terms of the triangle peak locations, $a_{i}$, rather than in terms of the network’s weights.
-The pretraining step of our algorithm then involves learning and updating these parameters.
-This is done by setting individual neuron weights in accordance with the procedures described in this section, and then backpropagating the gradients to update the triangle peak parameters.
+We begin this section by describing our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth.
+We focus on architecting the weights of a 4-neuron-wide, depth $d$ ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
+As alluded to in Section 2.1, the overall strategy of our reparameterization and initialization algorithm is to associate a nonsymmetric triangle function with each layer of the network.
+This yields a reparameterization of the ReLU network in terms of the locations of each triangle’s peak on $(0,1)$, $a_{i}$, rather than in terms of the network’s weights.
+Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions.
+Furthermore, this parameterization is trainable.
+By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions.
 Pseudocode for our entire algorithm is provided in Appendix A.1.
@@
 Each layer considered in isolation would compute these if directly fed the input signal.
-$T_{i}(x)$’s derivatives are the piecewise linear functions:
+$T_{i}(x)$’s derivatives are the piecewise constant functions:
 $T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
@@
 where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
-Figure 2: On the right is a network representation of a triangle function.
-The middle shows that triangle function as a hidden layer of a network.
+Figure 2: On the top left is a network representation of a triangle function.
+The top right shows that triangle function as a hidden layer of a network.
 The one-dimensional input and output of a triangle function is converted into shared weights.
-A full pretraining network is assembled on the left.
+A full pretraining network is assembled on the bottom.
 To encode these functions into the weights of a ReLU network, we begin with triangle functions (see the right subnetwork in Figure 2).
@@
 Figure 3: Each colored line shows the output signal of a neuron with respect to the input to the network. Colors match the corresponding neurons in Figure 2.
-Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep.
-We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead.
+Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$ shape, but this is unnecessarily deep and irregular.
+We can replicate the one-dimensional function composition in the hidden layers of our network (see Figure 2 top right).
 Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion.
 In this way, we can avoid having to use extra intermediate neurons.
-This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
+This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons possess identical slopes in regions where both neurons are active.
 There are two other neurons in the pretraining network’s layers, each with a specific role.
@@
 The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.
-We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially in order to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
+We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
+3.1 Pretraining and Overall Algorithm
+We imagine the process of optimizing neural network parameters as three phases: (1) reparameterization and initialization, where a network’s weights are initially set; (2) pretraining, where the network is trained under that reparameterization; and (3) training, where the network is trained using the network’s weights directly.
+Our pretraining algorithm acts as a preconditioner for or guide to the loss landscape, and when an optimizer is close to a minimum, the exponential expressivity of the network may no longer need to be guaranteed to ensure efficient neuron usage.
+Our pretraining algorithm is thus to simply use an optimizer like gradient descent to train the network under the reparameterization described above.
+Explicit formulas for updating the reparameterized weights, and pseudocode for our whole algorithm, are listed in Appendix A.1.
 3.2 Differentiable Model Output
@@
 For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
-For a broader mathematical discussion, as well as sufficient conditions for differentiability in the limit, see the appendix.
+Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3.
 Theorem 3.1.
@@
 4 One-Dimensional Experiments
-The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function.
-Accordingly, in this section, we implement our method to learn several convex one-dimensional curves.
-These nonlinear functions are known—so that we can measure error—and we remark and demonstrate results on more difficult functions in Section 5, Section 6, and Appendix A.10.
-The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
+Since our network construction applies directly to one dimensional convex functions, this will be the focus of this section (we cover more difficult functions in Sections 56 and A.10A.11) The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
 To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both.
-For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4).
-We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3).
-We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability.
-The intent of both comparisons is to see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape.
+We also train several permutations of our reparameterized networks, which validate that our pretraining and our differentibility constraints indeed facilitate smoother navigation of the loss landscape.
 Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
 4.1 Experimental Setup
-All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix for a brief analysis of different learning rates).
+All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix A.9 for a brief analysis of different learning rates).
 Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output.
 The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type.
-The networks unique to this paper share a common set of starting locations using the construction in Section 3, so that the effects of each training regimen are directly comparable.
-As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we focus for the moment on one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
+The networks using triangle based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4).
+In our tables, “pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization.
+“differentiability not enforced” is a network that initializes with our parameterization, but then trains the scales of its triangle waves as an extra degree of freedom (so that the final output is not necessarily differentiable).
+“differentiability enforced” corresponds to the main variant of our method as we have presented it thus far.
 The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave.
-The curves are chosen to capture a variety of convex one-dimensional functions.
 To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$.
@@
 We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible.
-The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.
+This set of experiments evaluates the predictive capacities of the networks on unseen data.
 4.2 Numerical Results
@@
 Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends.
-First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization.
+First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization (See also Figure 5).
 Even the networks that forgo pretraining benefit from initializing with many activation regions.
@@
 RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
-Figure 4: Comparison between standard Kaiming initialization/gradient descent and pretraining with differentiablity enforced.
-Using more linear regions allows the curve to better predict the test points.
+Figure 4: Standard Kaiming initialization/gradient descent vs. pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
 Our second set of results is shown in Table 3 and in Figure 4.
@@
 This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
-4.3 Gradient Descent does not Directly Optimize Efficiency
-Figure 5: Approximation produced with standard methodologies (shown before ReLU is applied). The neuron colors here are arbitrary; they do not correspond to Figure 2.
-Figure 5 shows the interior of a default network.
-The layers here are shown before applying ReLU.
-The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points.
-Examining the figure, the first two layers are wasted.
-No neuron’s activation pattern crosses $y=0$, so ReLU is never used.
-Layer 3 could be formed directly from the input signal.
-Deeper in the network, more neurons remain either strictly positive or negative.
-Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time.
-The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many.
-This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications.
-If a bend exists, gradient descent can reposition it.
-But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments.
-Therefore, bend-related information will be absent from its gradients.
-Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values.
-Occasionally, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
-4.4 Effect of Reparameterized Initialization vs Pretraining
-Figure 6: Pretraining (right) results in better structural retention in deeper layers.
-Figure 6 compares the top performing reparameterized models for $x^{3}$ with and without pretraining.
-We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure around layer 4 or 5, devolving into noisy patterns and resulting in higher errors.
-Pretraining maintains structure at greater depths.
-This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results.
-Gradient descent greedily abandons any such structure during its optimization in favor of models that can be worse in the long term.
-A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
-This phenomenon highlights the need for theoretical work to derive more expressive training controls, which can give better guidance for the optimization of functions of greater dimension or complexity.
+Figure 5: Neuron outputs of a default (Kaiming-initialized) network (left) versus a pretrained variant of our network (right).
+The reason our network succeeds in using so many linear regions, when a random one fails, is that the number of regions is not a “local” property of parameter space.
+The first 3 layers of the default network are not adding any new linear regions, and infinitesimal adjustments to the slopes or biases of the lines shown would not change that.
+In our network, the pretraining guides it to a region of parameter space where $2^{d}$ activation regions are generated.
+Since the region count is not a local property, we still have a neighborhood of parameter space we can optimize in and maintain efficient neuron use.
 5 Extension to Non-Convex Functions and Higher Dimensions
-Figure 7: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
-The results presented so far are limited to one-dimensional convex functions, but these are not limitations of our method.
+Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$.
+A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
+Notice though that the individual neuron outputs are much messier than in our convex case (see Figure 5) This is because naive gradient descent has all the weights in between the two pretrained component networks to learn unassisted.
+The results presented so far are limited to one-dimensional convex functions, but our method can easily be extended.
 Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 8 in the Appendix for a schematic).
@@
 We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising.
-In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world datasets and problems with up to eight input dimensions.
+In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world regression datasets with up to eight input dimensions, as well as image classification problems.
 We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1.
-This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957; 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions.
+This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions.
 But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees.
 Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
-Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks.
+Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks.
 Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect).
@@
 Figure 8: Extending our results into a larger network. ReLU should not be used in the layers between the blocks. Colors correspond to Figure 2.
-A.1 Initialization and Training Algorithm
+A.1 Initialization and Pretraining Algorithm
 The initialization step of our algorithm is to generate a vector $A=[a_{0},a_{1},...a_{n}]^{T}$, where each $a_{i}$ is randomized in $(0,1)$.
@@
 […]
-     A-Gradient $\leftarrow\text{Derivative}(\text{Network},A)$triangleright Backpropagate through weight setting
+     A-Gradient $\leftarrow\text{Derivative}(\text{Network},A)$Backpropagate through weight setting}
 […]
-     $A\leftarrow A-\epsilon\times\text{Gradient}$triangleright Update A, Not the network
+     $A\leftarrow A-\epsilon\times\text{Gradient}$Update A, Not the network}
   end while
@@
 The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
+Figure 9: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
 Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$.
@@
 This odd split around layer $i$ makes the proofs cleaner.
-Figure 10: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
 Figure 9 highlights some important properties about composing triangle functions.
@@
 We used 16 of our convex blocks, so all weight matrices are 64 by 64.
-Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 15 and 16 already show impressive results using our current approach.
+Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 14 and 15 already show impressive results using our current approach.
 In the case of the concrete problem (Figure 14), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced.
@@
 Further interesting optimizations of our approach (e.g., training another network to inform our algorithm when to switch parameterizations) are imagined as future work.
-A.11 Large-Scale Image Classification Examples
+A.11 Larger Image Classification Examples
 We include very preliminary demonstrations of how our method can increase performance on large-scale image classification tasks.
@@
 We evaluated two standard benchmarks: CIFAR-10 and ImageNet.
-In the case of CIFAR-10, we trained a small network of 3 convolutions followed by 2 fully connected layers.
+In the case of CIFAR-10, we trained a small network of three convolutions followed by two fully connected layers.
 To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added.
@@
 We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization.
-We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImagnNet.
-Figure 17: results for CIFAR-10 experiments.
-Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts; for instance, at epochs 3 and 4, which per the loss curves’ characteristics are where one would likely stop training.
-While the loss curves are similar in shape, we report that at epoch 4, for example, our training and validation losses are 25.6% and 5.1% lower than their standard network counterparts.
-Figure 18: In our ImageNet test we similarly see modest gains of our (“experimental”) network over a Kaiming-initialized (“ordinary”) network, yet due to using a smaller learning rate here ($0.00001$, vs.0.0010.0010.0010.001 as used with CIFAR-10), the loss curves are smoother, so we see our networks consistently outperforming the standard network 25% into epoch 1.
-We trained both networks for 2 epochs; at epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).
+We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImageNet.
+Note that in both experiments, we switch parameterizations in our method halfway through training.
+Figure 16: results for CIFAR-10 and ImageNet experiments.
+Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts.
+Interestingly, the results on ImageNet are much stronger (note the color labeling is the same for both images, our train and validation losses are the bottom two in ImageNet, whereas in Cifar-10 the two training losses are the lowest).
+This may be because the ImageNet experiments re-used the convolutional layers from VGG, rather than learning them on its own.
+It could also be in part due to larger layer sizes, since most activation functions are universal with infinite width, any detrimental effect of constraining them into convex shapes may be softened.
+In our ImageNet experiment we trained both networks for 2 epochs.
+Our method required a lower learning rate ($0.00001$) than used in previous experiments.
+At epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).
 These two examples show that even for large-scale image classification tasks—far away from the analytic 1D function tests considered in the main body—our methodology still has the potential to yield noticeable benefits, highlighting the promise of extending the present work to more general network architectures as a future research direction.

### v7

@@
 This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
-We further demonstrate how to extend our approach to multidimensional and non-convex functions, allowing it to replace the dense layers in other networks; preliminary improvements are shown for image classification on CIFAR-10 and ImageNet.
-Machine Learning, ICML
+We further demonstrate a preliminary extension of our construction to multidimensional and non-convex functions, allowing the technique to replace traditional dense layers in various architectures.
+ReLU networks, network initialization, linear regions, activation regions
+[…]
+mytheoremm
+[…]
 1 Introduction
@@
 The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] See Appendix A.2 for definitions of terms like linear regions, activation patterns, and activation regions.
-and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018).
+and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014; Serra et al., 2018).
 Hence deep architectures may outperform shallow ones.
@@
 Whether using a deep or shallow architecture, only the total number of neurons matters.
-They further observed that gradient descent has a difficult time creating new activation regions so that their bounds approximately held after training.
-As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see 5).
-Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
+They further observed that gradient descent has a difficult time creating new activation regions, and thus their bounds approximately held after training.
+As we will discuss later, the total number of linear regions is not a “local” property in parameter space that gradient descent can directly optimize (see Figure 5).
+Gradient descent is also prone to redundancy; for instance, Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
 The present work presents mathematical algorithms that avoid the limitations of random initialization. Our contributions include:
@@
   • Numerical results for one-dimensional test cases that yield orders-of-magnitude improvements in network performance (Section 4)
-  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Section 5; Appendices A.10A.11)
-While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and quickly outpacing such networks’ performance on more practical test cases such as ImageNet (Appendix A.11).
+  • Extensions to non-convex and multidimensional functions, and use of our method as a drop-in replacement for dense layers in arbitrary networks (Sections 5 and 6; Appendices B.2B.3)
+While an exponential increase in the expressiveness of a ReLU network does not necessarily imply an exponential increase in performance, one may intuitively expect a substantial benefit, and our results bear this out, yielding error values orders of magnitude lower than a traditionally-trained network of equal size for simple test problems (Section 4), and outpacing such networks’ performance during pretraining on more practical test cases such as ImageNet (Section 6).
 We emphasize that while the preliminary mathematical analysis in this work pertains to a specific ReLU construction, this construction can be repeated and combined in useful ways to obtain performance improvements on larger networks (Section 5).
@@
 Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
-The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)).
+The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014)).
 Our reparameterization strategy (Section 3) focuses on non-symmetric triangle waves.
@@
 Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon.
-In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
+In Section 5, we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
 Other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022), including spline-type methods (Eckle & Schmidt-Hieber, 2019).
@@
 Our work seeks to improve network initialization by making use of explicit theoretical constructs.
-This stands in sharp contrast the current standard approach, which treats neurons homogeneously.
+This stands in sharp contrast current standard approaches, which treat neurons homogeneously.
 Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010).
@@
 3 Initialization and Pretraining Construction
-We begin this section by describing our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth.
-We focus on architecting the weights of a 4-neuron-wide, depth $d$ ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
+In this section we describe our ReLU network reparameterization strategy that ensures an exponential number of linear regions with respect to depth.
+We focus on architecting the weights of a 4-neuron-wide, arbitrary depth ReLU network; while this is a constraint of the present manuscript, we will show that despite this, we can effectively apply our construction to a wide variety of examples.
 As alluded to in Section 2.1, the overall strategy of our reparameterization and initialization algorithm is to associate a nonsymmetric triangle function with each layer of the network.
@@
 To illustrate how we set weights, we first describe the mathematical functions that arise in our analysis. Triangle functions are defined as
-$T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases}$
+$T_{i}(x)=\begin{cases}\frac{x}{a_{i}}&0\leq x\leq a_{i}\\ 1-\frac{x-a_{i}}{1-a_{i}}&a_{i}\leq x\leq 1\end{cases},$
 where $10<a_{i}<$ 1.
-This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$.
-Each layer considered in isolation would compute these if directly fed the input signal.
-$T_{i}(x)$’s derivatives are the piecewise constant functions:
-$T^{\prime}_{i}(x)=\begin{cases}\frac{1}{a_{i}}&0<x<a_{i}\\ \frac{1}{1-a_{i}}&a_{i}<x<1\end{cases}$ (1)
-In a deep network, the layers feed into each other, composing their respective triangle functions:
-$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$ (2)
+This produces a triangular shape with a peak at $x=a_{i}$ and zeros at $x=0$ and $x=1$.
+These functions are implemented by two ReLU neurons in each layer.
+In a deep network, each layer composes its triangle function with the output of the layer before it.
+The result is that each layer computes the following triangular waveform:
+$W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x))).$ (1)
 These triangle waves will have $2^{i}$ linear regions, doubling with each layer.
-The output of the network in pretraining will be a weighted sum over the triangle waves formed at each layer.
-Assuming the network to be infinitely deep, we have
-$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$ (3)
+By allocating an extra neuron in each layer, our construction will take a weighted sum over each layer’s triangle wave, using them as a deep basis to build the output.
+In the infinite depth limit, the network output takes the form:
+$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x),$ (2)
 where $s_{i}$ are scaling coefficients on each of the composed triangular waveforms $W_{i}$.
+$F(x)$ would be a parabola if we exactly followed the related work.
+$F(x)$ could also end up being a fractal curve if the $s_{i}$ are chosen to be too large.
+Choosing $s_{i}$ carefully to regularize $F(x)$ is a major part of our construction.
 Figure 2: On the top left is a network representation of a triangle function.
@@
 A full pretraining network is assembled on the bottom.
-To encode these functions into the weights of a ReLU network, we begin with triangle functions (see the right subnetwork in Figure 2).
-Its maximum output is $1$ at the peak location $a\in(0,1)$.
-Neuron $t_{1}$ simply preserves the input signal.
-Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$.
-Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal.
-The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope.
-When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously.
-Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves.
-In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions (see Figure 3).
+Now we will discuss how to encode these functions into the weights of a ReLU network, beginning with a network for a single triangle function (see the upper left diagram in Figure 2).
+Each triangle function has two nonzero linear pieces, requiring two ReLU neurons.
+For simplicity, we will set both neurons’ input weights to $1$.
+We will arrange for neuron $t_{1}$ to be active for all inputs, and since the maximum output should be $1$ at the peak location $a\in(0,1)$, neuron $t_{1}$ will have an outgoing weight of $1/a$.
+$x=a$ is where the slope should change, so neuron $t_{2}$ will be biased by $-a$ to suppress it for $x<a$.
+Since both neurons are active for $x>a$, neuron $t_{2}$ will need to be weighted by $-(1/a+1/(1-a))=-1/(a-a^{2})$; this weight will both suppress neuron $t_{1}$ and also ensure the second leg of the triangle function reaches 00 at $x=1$.
+This is now the first network shown in Figure 2.
+To compose the triangle waves $W_{i}(x)$, copies of this network can be stacked together depth-wise, lining up the “in” and “out” neurons of successive networks.
+This will have a non-uniform width ($1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$), but since the input weights of the hidden $t_{1}$ and $t_{2}$ neurons are both $1$, it can be simplified into the form shown on the top right of Figure 2.
+The weights $1/a$ and $-1/(a-a^{2})$ give the correct way to combine the $t_{1}$ and $t_{2}$ neurons in order to obtain a triangle function, so every hidden neuron will combine the previous $t_{1}$ and $t_{2}$ neuron in this proportion.
+Using identical weights makes each layer act as if it is composing functions in one dimension, rather than in a dimension equal to the number of neurons.
+The goal of the network is not to output triangle waves; instead, the function we seek to approximate is $F(x)$, defined earlier to be a weighted sum of the triangle waves from each layer.
+We add an extra neuron to each layer, labeled “sum” in Figure 2, to compute this.
+These are ordinary neurons, but thanks to their specific weights, they will act similarly to a residual connection (He et al., 2016), allowing the hidden features in the network to be used in the output.
+Each sum neuron adds the previous layer’s triangle wave to the previous sum neuron, thereby iteratively updating an approximation to $F(x)$.
+According to Theorem 3, each triangle wave added to the sum must have an amplitude exponentially smaller than the last.
+Exponentially small weights would pose conditioning issues in optimization, so rather than giving the sum neuron an exponentially small weight, we use the ratio of successive scaling coefficients $S_{i}=s_{i}/s_{i-1}$ to iteratively decay the amplitude of the $t_{1}$ and $t_{2}$ neurons.
+But because $t_{2}$ now has an exponentially small amplitude, its bias must be exponentially small to function properly.
+Rather than implementing an exponentially small bias, we make the bias into a neuron.
+This is rather unconventional, but by using the connection to the previous bias neuron, the bias can decay iteratively, and $t_{2}$ neurons can use their weight to the bias neuron to implement their bias.
+In matrix form, our hidden layers are:
+$\begin{bmatrix}1&\pm[S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})]&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times\begin{bmatrix}\text{sum}\\ t_{1}\\ t_{2}\\ \text{bias}\end{bmatrix}.$
+Here, rows correspond to neurons, and the sum neuron can either add or subtract the triangle waves depending on whether the target function is convex or concave.
 Figure 3: Each colored line shows the output signal of a neuron with respect to the input to the network. Colors match the corresponding neurons in Figure 2.
-Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow 1$ shape, but this is unnecessarily deep and irregular.
-We can replicate the one-dimensional function composition in the hidden layers of our network (see Figure 2 top right).
-Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion.
-In this way, we can avoid having to use extra intermediate neurons.
-This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons possess identical slopes in regions where both neurons are active.
-There are two other neurons in the pretraining network’s layers, each with a specific role.
-The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer.
-This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions).
-If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients.
-Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues.
-Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer.
-A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values.
-Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias.
-The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.
-We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.
 3.1 Pretraining and Overall Algorithm
@@
 Given the rapid oscillations of the triangle waves formed at each layer, the network will output a fractal with many choices of scaling parameters.
-This would be poorly predictive of unseen data points generated from a smooth curve.
-Our main mathematical result addresses this issue by forcing differentiability of the network output.
-We do so by making the sum of the derivatives of the triangle waves converge.
-For this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
-Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3.
+This would be poorly predictive of unseen data points generated by a smooth curve.
+Rather than turning to a general regularizer like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), we use a regularization scheme that arises from the form of our functions.
+When we require that the network output approach a differentiable function in the infinite depth limit, an elegant relationship arises where the peak locations of the triangle waves uniquely determine the scales with which to sum them.
 Theorem 3.1.
 $F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$ and is continuous on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
-$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$ (4)
+$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}.$ (3)
+Mathematical analysis and sufficient conditions for differentiability in the limit are in Appendix A.3.
+In Section 4, we will see that following this regularization allows the pretraining phase to produce better initial solutions.
 4 One-Dimensional Experiments
-Since our network construction applies directly to one dimensional convex functions, this will be the focus of this section (we cover more difficult functions in Sections 56 and A.10A.11) The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
-To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both.
+Since our network construction applies directly to one-dimensional convex functions, we focus on such functions in this section (we cover more difficult functions in Sections 56 and Appendices B.2B.3).
+The aim of these experiments is twofold: (1) we would like to determine the most effective function representations possible from such a small network, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data.
+To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution (Singh & Sreejith, 2021), and produce errors that are orders of magnitude lower than both.
 We also train several permutations of our reparameterized networks, which validate that our pretraining and our differentibility constraints indeed facilitate smoother navigation of the loss landscape.
@@
 4.1 Experimental Setup
-All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix A.9 for a brief analysis of different learning rates).
+All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix B.1 for a brief analysis of different learning rates).
 Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output.
 The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type.
-The networks using triangle based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4).
-In our tables, “pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization.
-“differentiability not enforced” is a network that initializes with our parameterization, but then trains the scales of its triangle waves as an extra degree of freedom (so that the final output is not necessarily differentiable).
-“differentiability enforced” corresponds to the main variant of our method as we have presented it thus far.
+To compare how well each setup navigates the loss landscape, the networks using triangle-based parameterizations share a common set of starting locations where the triangle peaks and scaling parameters are chosen according to Theorem 3.1.
+The “unregularized” network has extra degrees of freedom to choose the scaling coefficients after initialization (so that differentiability of the output is not enforced during pretraining).
+“pretraining skipped” is a network that initializes with our method, but only trains in the standard parameterization.
 The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave.
-To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$.
-For the other approximations, the waves are subtracted.
-This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$.
+To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$, while for $x^{3}$ and $x^{11}$ the waves are subtracted.
+When the waves are subtracted, the first scaling factor has to be changed to $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$.
 The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves.
-This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network.
+This is chosen to be very dense deliberately, to try to evoke the most accurate function interpolations from each network.
 We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments.
@@
 Pretraining Skipped $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
-Differentiability Not Enforced $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
-Differentiability Enforced $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
+Pretraining (Unregularized) $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
+Pretraining (with Thm. 3.1) $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
 Table 2: Minimum and mean (30 samples) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
@@
 Pretraining Skipped $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
-Differentiability Not Enforced $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
-Differentiability Enforced $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
+Pretraining (Unregularized) $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
+Pretraining (with Thm. 3.1) $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
 Table 3: Minimum errors on unseen points from training on sparse data.
@@
 Default Network (Kaiming) $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
-Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
+Pretraining (with Thm. 3.1) $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
 Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends.
-First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization (See also Figure 5).
+First, the worst performing networks are the “default networks” that rely on randomized (Kaiming) initialization (see also Figure 5).
 Even the networks that forgo pretraining benefit from initializing with many activation regions.
@@
 The last trend to observe is the poor average performance of default networks.
-In a typical run of these experiments, around half of the default networks collapse.
+In a typical run of these experiments, around half of the default networks collapse from the dying ReLU phenomenon.
 RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
+This is why we have focused on presenting the minimum error over all the trials.
 Figure 4: Standard Kaiming initialization/gradient descent vs. pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
@@
 Here the most important impact of utilizing exponentially many linear regions is demonstrated.
-Not only can more accurate representations of training data be learned, but more linear regions allow the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks.
+Not only can more accurate representations of training data be learned, but maintaining more linear regions allows the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks.
 This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
 Figure 5: Neuron outputs of a default (Kaiming-initialized) network (left) versus a pretrained variant of our network (right).
-The reason our network succeeds in using so many linear regions, when a random one fails, is that the number of regions is not a “local” property of parameter space.
-The first 3 layers of the default network are not adding any new linear regions, and infinitesimal adjustments to the slopes or biases of the lines shown would not change that.
-In our network, the pretraining guides it to a region of parameter space where $2^{d}$ activation regions are generated.
-Since the region count is not a local property, we still have a neighborhood of parameter space we can optimize in and maintain efficient neuron use.
+Notice that the first two layers of the default network introduce no linear regions - none of the lines cross zero.
+Any infinitesimal adjustment to the slopes or biases of the lines would not make such an intersection occur.
+Therefore, the number of linear regions generated by the network cannot be a local property, and we can expect gradient-based optimization to struggle at maximizing the linear region count.
+Our method uses this non-locality to our advantage.
+The pretraining phase finds a low-loss solution where $2^{d}$ linear regions are generated, which guarantees a neighborhood of parameter space where $2^{d}$ regions can be maintained while training in the standard parameterization.
 5 Extension to Non-Convex Functions and Higher Dimensions
-Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$.
-A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
-Notice though that the individual neuron outputs are much messier than in our convex case (see Figure 5) This is because naive gradient descent has all the weights in between the two pretrained component networks to learn unassisted.
+Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $5.52\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
 The results presented so far are limited to one-dimensional convex functions, but our method can easily be extended.
-Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 8 in the Appendix for a schematic).
-For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs.
-Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
-The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al.
-(2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions).
-We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising.
-In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world regression datasets with up to eight input dimensions, as well as image classification problems.
-We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1.
-This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions.
-But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees.
-Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
-Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks.
-Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect).
-Two slight modifications are made to each of the component networks to better interact with ReLU.
-First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU.
-Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
+Since our constructed networks are one-dimensional, they can be used as the activation functions of a larger network (see Figure 9 in the Appendix for a schematic).
+Our experimental results show that even using just two copies of our network in this manner can go a long way towards increasing the expressive ability of the technique.
+For example, two of our networks could be used to build a two-dimensional convex function.
+This can be done by using one network to build a convex function oriented in the $x$ direction, and the other in the $y$ direction.
+A positive linear combination of their outputs then yields a two-dimensional convex function.
+Similarly, a non-convex function can be realized by taking a difference of the outputs of two of our networks.
+Of course, this pattern can be generalized to an arbitrary-width layer using $n$ of our networks.
+Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional convex function $z=r^{3}$ on $[-1,1]$, each by using two of our constructed networks with 5 hidden layers.
 For both $x^{3}-x$ and $r^{3}$, we substantially outperform standard initialization and training procedures.
 Especially striking is the increased number of linear regions in Figure 7.
-Visible in Figure 6 is the gradient descent-induced destructuring discussed in Section 4.2.
-This likely occurs due to the block diagonal structure of the larger weight matrices, which results in many zeros for the second pass of gradient descent to fill in unassisted.
-This could be addressed either by limiting training only to the block diagonal entries, or by deriving a more advanced pretraining method that makes use of the full matrix.
-Note that since $x^{3}-x$ is twice differentiable we can write it as $(x^{3}-x+3x^{2})-3x^{2}$, a difference of functions convex on $[-1,1]$ obtained by finding a parabola based on trying to make the derivative of $x^{3}-x$ monotone.
-Note that this is not a unique decomposition.
+Like the previous experiments, these networks were trained using Adam for 1000 epochs at a learning rate of 0.001, and the minimum over 30 trainings was taken.
+The parameterization switch was made halfway through training.
+These experiments use a slightly improved pretraining construction (see Appendix A.1): the input domain is modified to be $[-1,1]$, and the network is adjusted so that the $t_{1}$ and $t_{2}$ neurons compose “v” shapes rather than triangles, which allows the network to extrapolate better outside $[-1,1]$.
+With multiple copies of the network, we implement the calculation of the $d$-th hidden layer as one large combined matrix.
+Since the hidden neurons of each subnetwork do not connect, the combined matrix takes on a sparse block diagonal structure with a block size of 4.
+When we switch parameterizations, we release the block diagonal constraint and allow gradient descent to fill in the zeroed weights.
+We make a few notes on our preliminary extended technique.
+First, the class of non-convex functions that can be represented as a difference of convex functions is broad.
+Zhang et al.
+(2018) gives an iterative process to decompose an arbitrary ReLU network into a difference of piecewise linear convex functions.
+Additionally, all functions with bounded second derivatives (or bounded eigenvalues of their Hessian matrix) can be expressed as a difference of two convex functions.
+Finally, we note that these types of decompositions are not necessarily unique; for example, while $x^{3}-x$ can be expressed as $(x^{3}-x+3x^{2})-3x^{2}$, $3x^{2}$ is merely the minimal parabola needed to make the second derivative positive on $(-1,1)$, and larger coefficients could also be used.
+Even though our original technique is not naturally designed for multivariate functions and has to be extended by incorporating it as an activation function in a larger network, we note that doing so has some theoretical backing.
+The multiplication gate and polynomial interpolation network discussed in the related work (Section 2.1) rely on linear combinations of squared terms.
+Since our networks can produce $x^{2}$, an arbitrary width and depth arrangement of our networks would retain the ability to perform polynomial interpolation.
+Secondly, the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959) gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure that is built using only addition and arbitrary continuous one-dimensional activation functions.
+Essentially, it states that a technique that only works for univariate functions can still be sufficient for representing multivariate functions when deployed in a larger network.
+The recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024) take their inspiration from this theorem, extending their networks to arbitrary width and depth, but restricting the activations to splines.
+Even though this does not strictly follow the theorem, they still find empirical success.
+Similarly, in the case of this work, the functions that the triangular parameterization can directly represent are not dense, as they require a certain degree of self-similarity.
+Nonetheless, when assembled into a larger network, this may not pose a serious problem.
 Figure 7: Approximations of $z=\sqrt{x^{2}+y^{2}}^{3}$ using an $8\times 5$ regular network (left) and a union of two of our pretrained components (right).
 Losses are $1.5\times 10^{-4}$ and $3.5\times 10^{-6}$, respectively, demonstrating a nearly two orders of magnitude improvement using our techniques.
+6 Preliminary Image Classification Results
+Since our extended technique can be used to replace any dense layers in arbitrary architectures, we can use it in contexts beyond regression.
+In our final experiment, we replace the densely connected classifier of VGG-16 (Simonyan & Zisserman, 2014), and then retrain the network on ImageNet (Deng et al., 2009).
+We initialize with PyTorch’s pretrained convolutional weights both to speed up training and to better isolate the effects of modification to the dense layers.
+Since VGG-16’s dense layers have a width of 4096, we decided to leave the matrices corresponding to our subnetworks in block diagonal format, rather than allowing them to become dense.
+This way, our technique only adds a negligible amount of extra operations (roughly $0.5\%$) and does not leave the majority of weights to be filled in by unassisted gradient descent.
+Figure 8: Retraining the classifier of a VGG-16 model on ImageNet.
+The ordinary network uses a learning rate of $1\times 10^{-4}$, and our experimental network uses $1\times 10^{-5}$.
+At epoch 10, both learning rates decay by a factor of 10, and we switch the parameterization in the experimental network.
+Both networks retrain slightly higher than PyTorch’s original reported accuracy of $73.3\%$.
+Figure 8 shows that our network has an advantage early in training, but the fully trained networks achieve comparable accuracy.
+This is perhaps unsurprising: ordinarily initialized and trained networks are already extremely good at classification problems despite the pathological properties of random initialization.
+This result suggests that highly precise representations of decision boundaries may not be critical to performance for such classification problems.
+In support of this, we noticed that it did not seem to matter if we switched parameterizations with our technique, even though doing so was critical to our performance on the 1D regression problems.
+A practical-scale regression task may be needed for our method to display a significant advantage; we plan such examples as future work.
 7 Concluding Remarks
@@
 This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks.
-Although we proposed one preliminary strategy for extending our network for higher-dimensional functions, this strategy has limitations, such as leaving many weights to be filled in by gradient descent (which it generally does not do effectively).
-We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture.
-We note that there are other metrics for evaluating a neural network’s expressivity (e.g., Raghu et al.
-(2017)), which would be interesting to measure for our networks in future work.
+We anticipate the continued development of theoretical improvements to this work.
+Of particular importance, we think, is either finding a dense set of one-dimensional functions we can represent efficiently with deep networks, or finding a method that more naturally represents multidimensional functions.
+While works like the KAN show it is possible to perform well without those things, we believe them to be important.
+The fact that our current technique is limited to creating sparse block diagonal matrices is indicative of a need to determine how to use all the weights of dense layers.
+Another interesting avenue for increasing the technique’s expressiveness could be to consider the generation of fractals.
+We disallow such behavior by using Theorem 3.1 to enforce differentiability in the hopes of regularization, but there may be more sophisticated approaches to function modeling that can generate fractals in a controlled manner.
+Given that the technique in its current form is flexible enough to go into any architecture, it is also important to explore possible use cases.
+We have shown that it will likely not lead to significant improvements on classification tasks, but the technique may shine when given an appropriate utility-scale regression task.
 Overall, we are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU and other classes of neural networks.
-Appendix A Appendix
+Impact Statement
+This paper presents work whose goal is to enable more efficient neural networks.
+Future advances along this line could enable the use of much smaller networks in practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
+Acknowledgments
+M.M.
+would like to thank Xander Neuwirth for insightful conversations.
+D.H.
+was supported in part by the U.S.
+National Science Foundation under award number 2442853.
+This work was authored in part by the National Renewable Energy Laboratory (NREL), operated by Alliance for Sustainable Energy, LLC, for the U.S.
+Department of Energy (DOE) under Contract No.
+DE-AC36-08GO28308.
+This work was supported by the Laboratory Directed Research and Development (LDRD) Program at NREL.
+The views expressed in the article do not necessarily represent the views of the DOE or the U.S.
+Government.
+The U.S.
+Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S.
+Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S.
+Government purposes.
+Appendix A Algorithm and Theory
 Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks. Colors correspond to Figure 2.
@@
 The initialization step of our algorithm is to generate a vector $A=[a_{0},a_{1},...a_{n}]^{T}$, where each $a_{i}$ is randomized in $(0,1)$.
-Given this $A$, the pretraining step of our algorithm sets the weights of the input ($I$), hidden ($H_{i}$, $11\leq i\leq n-$ 1), and output ($O$) layers of the network as follows:
+Given this $A$, the pretraining step of our algorithm (in the experiments in Section 4) sets the weights of the input ($I$), hidden ($H_{i}$, $11\leq i\leq n-$ 1), and output ($O$) layers of the network as follows:
 $I(x)=\begin{bmatrix}x\\ x\\ x\\ 0\end{bmatrix}+\begin{bmatrix}0\\ 0\\ -a_{0}\\ 1\end{bmatrix}$
-$H_{i}(x)=\begin{bmatrix}1&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times ReLU(H_{i-1}(x))$
-$O(x)=\begin{bmatrix}1&S_{n}/a_{n}&-S_{n}/(a_{n}-a_{n}^{2})&0\end{bmatrix}\times ReLU(H_{n-1}(x))$
-where $S_{i}$ can either be chosen independently, or chosen based on $A$.
+$H_{i}(x)=\begin{bmatrix}1&\pm[S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})]&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times ReLU(H_{i-1}(x))$
+$O(x)=\begin{bmatrix}1&S_{n}/a_{n}&-S_{n}/(a_{n}-a_{n}^{2})&0\end{bmatrix}\times ReLU(H_{n-1}(x)),$
+where $S_{i}$ can either be chosen independently or chosen based on $A$.
 In the latter case, Equation 3 gives $S_{i}=s_{i}/s_{i-1}=(1-a_{i})a_{i+1}$.
-This assignment of $I$, $H$, and $O$ is used in each iteration of the pretraining algorithm.
+For this version, one must decide whether to add or subtract triangle waves from the sum neuron based on whether the target function is convex or concave, respectively.
+The version of the algorithm used in Section 5 onward does not have this stipulation and always follows Equation 3.
+Additionally, it expands the input domain to $[-1,1]$ and flips the triangles upside down to compose “v”-shaped functions instead:
+$I(x)=\begin{bmatrix}0\\ \frac{-0.5}{a_{0}}\\ \frac{0.5}{1-a_{0}}\\ 0\end{bmatrix}x+\begin{bmatrix}1\\ 1-\frac{0.5}{a_{0}}\\ 1-\frac{0.5}{1-a_{0}}\\ 1\end{bmatrix}$
+$H_{i}(x)=\begin{bmatrix}1&1&1&-1\\ 0&a_{i+1}-1&a_{i+1}-1&a_{i}(1-a_{i+1})\\ 0&\frac{a_{i}(1-a_{i+1})}{1-a_{i}}&\frac{a_{i}(1-a_{i+1})}{1-a_{i}}&\frac{-a_{i}^{2}(1-a_{i+1})}{1-a_{i}}\\ 0&0&0&a_{i}(1-a_{i+1})\end{bmatrix}\times ReLU(H_{i-1}(x))$
+$O(x)=\begin{bmatrix}1&1&1&-1\end{bmatrix}\times ReLU(H_{n-1}(x)).$
+This assignment of $I$, $H$, and $O$ is used in each iteration of the pretraining algorithm to set the weights of each layer before the forward pass.
+The backward pass can then propagate the gradients back through this weight setting procedure to update the vector $A$ containing the triangle peaks.
+After pretraining, the weights can then be initialized once more based on the learned vector $A$, and then updated directly using regular gradient descent in a second round of optimization.
+Full pseudocode is listed in Algorithm 1.
 Algorithm 1 Initialization and Pretraining
@@
   end while
-The network weights can then be initialized once more based on the learned vector $A$, to then update the weights via regular gradient descent.
-In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs.
-Full pseudocode is listed in Algorithm 1.
 A.2 Terminology
 Before presenting further formal mathematical details of our method, we first briefly review a few pieces of basic ReLU network terminology used in the paper.
-The reader is referred to the Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
+The reader is referred to Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
 An activation pattern is a boolean mask that tracks which neurons in a network have their output zeroed by ReLU activations.
@@
 For instance, Rolnick & Kording (2020) leverages theoretical works like those cited in our related work discussion for an exciting privacy application: it is proven that a ReLU network’s output can often be provably used to reverse engineer the architecture of a ReLU network, up to isomorphism.
-A.3 Mathematical Results
+A.3 Necessary Conditions for Differentiability
 For convenience, we first restate the functions defined in the main body of the paper.
@@
 $W_{i}(x)=\bigcirc_{j=0}^{i}T_{j}(x)=T_{i}(T_{i-1}(...T_{0}(x)))$
-$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$
-The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$.
-We begin by assuming that
-$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$
+$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x).$
+The goal of this section is to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$. We begin by assuming that
+$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x).$
 We will see that the resulting choice of $s_{i}$ ensures uniform convergence of the derivative terms, so that the derivative of the infinite sum is indeed infinite sum of the derivatives.
-Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right respectively.
+Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right, respectively.
 The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
@@
 To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$.
-Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition.
+Each triangle function has two distinct slopes, $1/a_{i}$ and $-1/(1-a_{i})$, which are dilated by the chain rule during the composition.
 On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected.
@@
 Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$.
-Either way, it’s important to note that the slopes on each side of a bend scale identically during each composition.
+Either way, it is important to note that the slopes on each side of a bend scale identically during each composition.
 Before we begin reasoning about $F^{\prime}(x)$, it can simplify the analysis to only consider the derivative of the error function $E^{\prime}(x)$.
@@
 $T_{i+1}$ has two linear segments, giving two slope possibilities to multiply by.
-The correct one to choose is $-1/(1-a_{i+1})$ because it’s ‘active’ around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$).
+The correct one to choose is $-1/(1-a_{i+1})$ because it is “active” around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$).
 This gives $W^{\prime}_{i+1}(x_{k})=-R_{x}\frac{s_{i+1}}{(1-a_{i+1})}$.
@@
 Summing up all the terms with the coefficients $s_{i}$, and factoring out $R_{x}$ will yield the desired formula.
-Note that this same derivation applies to the left sided derivatives as well because the ‘active’ slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right.
+Note that this same derivation applies to the left sided derivatives as well because the “active” slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right.
 The initial slope constant $L_{x}$ will just be different.
@@
 For all $x\in P_{i}$:
-$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (6)
+$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x).$ (5)
 Proof.
@@
 Once they are both valleys, they will be treated the same by the remaining compositions, so the sizes of their remaining discontinuities will need to be proportional.
-Theorem (3.1).
-$F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
-$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$
+{mytheorem}3.1 $F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
+$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}.$
 Proof.
@@
 allows for a substitution to eliminate the infinite sum
-$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}$
+$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}.$
 Collecting all the terms gives
-$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}}$
+$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}},$
 which simplifies to the desired result. ∎
 Figure 11: The derivatives of the first few stages of approximation.
-Notice that each time a constant segment “splits”, the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it’s from)
+Notice that each time a constant segment “splits,” the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it originates from).
 A.4 Sufficiency for differentiability
@@
 Since the ratio of two successive scaling terms is $(1-a_{i+1})a_{i+2}$, factors of both $a_{i}$ and $1-a_{i}$ are present in $s_{i}$.
-This has a few implications, firstly if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth.
+This has a few implications.
+Firstly, if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth.
 Secondly, having both $a_{i}$ and $1-a_{i}$ will cancel whichever slope multiplier $T_{i}$ contributes to $W_{i}$ at each point $x$, leaving behind the other term, which is less than 1.
@@
 Proof.
-We begin by considering Equation 5 for layer $i$ (it equals 0 by Theorem A.3).
-$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
+We begin by considering Equation 4 for layer $i$ (it equals 0 by Theorem A.3).:
+$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$
 We will prove our result by substituting Equation 3 into this formula, and then verifying that the resulting equation is valid.
@@
 Equation 3 gives a recurrence relation.
-Converting it to an non-recursive representation we have:
-$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (7)
+Converting it to an non-recursive representation, we have:
+$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right).$ (6)
 When we substitute this into Equation 4, three things happen: each term is divisible by $s_{i}$ so $s_{i}$ cancels out, every factor in the product except the last cancels, and $1-a_{i+1}$ cancels.
 This leaves
-$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (8)
+$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}).$ (7)
 This equation has a meaningful interpretation that is important to the argument.
@@
 The following argument shows that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining discontinuity, guaranteeing its disappearance in the limit.
-Inductively we can show:
-$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (9)
-In words this means that as the first term appearing on the right in Equation 8 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
+Inductively, we can show:
+$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m}).$ (8)
+In words, this means that as the first term appearing on the right in Equation 7 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side.
 As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$.
-Assuming the above equation holds for all previous values of $j$
-$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
+Assuming the above equation holds for all previous values of $j$,
+$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m})),$
 using the inductive hypothesis to make the substitution
-$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
+$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m})).$
 Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded by the exponentially decaying series $(1-c)^{n}$, which approaches zero. ∎
@@
 by applying Equation 3 twice, we have
-$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
-To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
+$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}.$
+To maximize $s_{i+2}$, we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
 Since each $W_{i}$ takes values between 0 and 1, its contribution to $F$ is bounded by $s_{i}$.
@@
 $x$ is a peak location of $W_{i}$, so the right side slope is negative and the left is positive.
-Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})/L$ and $r_{1}=x+(1-a_{i+1})/R$.
+Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side gives $l_{1}=x-(1-a_{i+1})/L$ and $r_{1}=x+(1-a_{i+1})/R$.
 On each subsequent iteration $i+n$ ($n\geq 2$), $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$.
-Since $x$ is a valley point the right slope is positive and the left is negative.
+Since $x$ is a valley point, the right slope is positive and the left is negative.
 The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $W_{i+n}$ ranges from 0 to 1 over these spans.
-Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$.
+Solving for the new peaks again gives $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$.
 The resulting non-recursive formulas are:
-$x-l_{n}=\frac{1-a_{i+1}}{L}\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{1-a_{i+1}}{R}\prod_{m=2}^{n}a_{i+m}$ (10)
+$x-l_{n}=\frac{1-a_{i+1}}{L}\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{1-a_{i+1}}{R}\prod_{m=2}^{n}a_{i+m}.$ (9)
 The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
@@
 Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 8 with the initial discontinuity size set to $Rs_{i}$ rather than 1.
-Focusing on the right hand side we get:
-$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=R*s_{i}\prod_{m=2}^{n}(1-a_{i+m})$
+Focusing on the right hand side, we get:
+$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=R*s_{i}\prod_{m=2}^{n}(1-a_{i+m}).$
 Taking $\Delta y/\Delta x$ gives a series:
-$\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
-The issue which arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$.
-The only way for these formulas to agree then is for $R^{2}=L^{2}$ which implies $a_{i}=1-a_{i}=0.5$.
+$\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}.$
+The issue that arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$.
+The only way for these formulas to agree, then, is for $R^{2}=L^{2}$, which implies $a_{i}=1-a_{i}=0.5$.
 Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$).
@@
 A.7 Monotonicity and Continuity of Derivatives
-Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure10).
+Each of the $f^{\prime}_{i}$ are composed of constant value segments.
+We will show that those values are monotonically decreasing (this can be seen in Figure11).
 This can extend into the limit to show that $F^{\prime}$ is monotone decreasing and that $F$ is concave.
@@
 Proof.
-To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i+1}[2^{i}]]$ for $0\leq n\leq 2^{i}$, which tracks the values of $F^{\prime}$ at the $i$th set of valley points.
+To establish this, we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i+1}[2^{i}]]$ for $0\leq n\leq 2^{i}$, which tracks the values of $F^{\prime}$ at the $i$th set of valley points.
 All but the first and last points will have been peaks at some point in their history, so Equation 5 gives the value of those derivatives as $f^{\prime}_{i}$.
@@
 Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave.
-Which then be used to prove that their limit $F$ is also concave.
-As the base case $f_{0}$ is a line with derivative 0, $V_{0}$ contains its two endpoints.
-$Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore both the points in $Y^{\prime}$ are in descending sorted order.
+This can then be used to prove that their limit $F$ is also concave.
+As the base case, $f_{0}$ is a line with derivative 0, and $V_{0}$ contains its two endpoints.
+$Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes.
+Therefore, both the points in $Y^{\prime}$ are in descending sorted order.
 The second part of the invariant is true since 0 is in between those values.
-Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value, which we know from the second inductive hypothesis is in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$.
+Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$.
+This entire interval is between two valley points, so $f^{\prime}_{i}$ (which has not added $W_{i}$ yet) is some constant value, which we know from the second inductive hypothesis is in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$.
 The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$.
@@
 Then there exist points $a$, $b$, and $c$ such that $F(b)$ lies strictly below the line connecting the points $(a,F(a))$ and $(c,F(c))$.
-Lets say it’s below the line by an amount $\epsilon$.
+Let us imagine it is below the line by an amount $\epsilon$.
 Since at each point $f_{i}$ converges to $F$, we can find $i_{a}$ such that $f_{ia}(a)-F(a)<\epsilon/2$, etc…, we take $i=\max(i_{a},i_{b},i_{c})$.
@@
 ∎
-Lastly we briefly sketch out why $F^{\prime}$ is continuous.
+Lastly, we briefly sketch out why $F^{\prime}$ is continuous.
 It relies on some of our earlier reasoning.
@@
 Counting the number of cells, which are convex connected components of space separated by the hyperplanes, is equivalent to determining the number of activation regions produced by the shallow ReLU network.
-A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc… General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network).
+A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc.
+General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network).
 Zaslavsky’s theorem states that
@@
 When in general configuration, the number of cells ($R\binom{n}{D}$) created by a $D$-dimensional arrangement of $n$ hyperplanes is
-$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}$
-To compute this sum, one would go down to the row of Pascal’s triangle corresponding to the number of planes $n$, and sum the first $D$ entries (see Figure 12 for a visual example).
-It’s a known fact that the sequence of the $d$th entry for each row grows to the $d$th power.
-So as the number of hyperplanes tends towards infinity, this sum is dominated by the $D$th term, which grows to the power $D$.
-In other words, the number of linear regions formed by a single hidden layer ReLU network grows exponentially with respect to its input dimension.
-This property partly explains why neural networks are successful at learning high dimensional functions, even when other classical methods become intractable.
+$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}.$
+[…]
 Figure 13: 3 lines dividing 2-dimensional space creates 7 cells.
 This is calculated by going down to row 3 (for the number of lines) and summing the numbers at indices 0, 1, and 2 (for the dimension).
-Also note that there is one 2-D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
-Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e.
-points, lines, planes, etc…) are given by the entries in Pascal’s triangle.
-So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower dimensional objects, or that each lower dimensional object could be ‘assigned’ a cell.
-There’s a striking similarity in the intuitions behind the recurrence that generates pascal’s triangle, and the calculation of the numbers of lower dimensional objects created from a hyperplane arrangement.
+Also note that there is one 2D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
+Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e., points, lines, planes, etc.) are given by the entries in Pascal’s triangle.
+So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower-dimensional objects, or that each lower-dimensional object could be “assigned” a cell.
+There is a striking similarity in the intuitions behind the recurrence that generates Pascal’s triangle, and the calculation of the numbers of lower-dimensional objects created from a hyperplane arrangement.
 In the case of Pascal’s triangle, if one wished to know how to choose 3 items from 6, one could reason that if item 6 is included, 2 items would be chosen from the remaining 5, and if item 6 is excluded, 3 would be chosen from the remaining 5, and therefore $\binom{6}{3}=\binom{5}{2}+\binom{5}{3}$.
 Now consider the case of 3 planes partitioning 3-space, and the addition of a 4th plane.
-The total number of lower dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points).
-One could also picture that the new 4th plane ‘cuts’ the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 12), and count the 2-cells that appear.
+The total number of lower-dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points).
+One could also picture that the new 4th plane “cuts” the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 13), and count the 2-cells that appear.
 Either reasoning leads to $R\binom{n}{D}=R\binom{n-1}{D}+R\binom{n-1}{D-1}$.
@@
 Inducting over Pascal’s triangle (with a base case of $R\binom{n,m}{1}=1+nm$) would give the non-recursive formula of $R\binom{n,m}{D}=\sum_{d=0}^{D}m^{d}\binom{n}{d}$.
-This means that out architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic trade off once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations).
+This means that our architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic tradeoff once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations).
 Of course, the effect on accuracy will all depend on how well these regions can be utilized.
-A.9 Learning Rates
+Appendix B Additional Experiments
+B.1 Learning Rates
 Figure 14: Learning $y=x^{3}$ at various learning rates for 1000 epochs.
@@
 For low learning rates (below $10^{-5}$), the steps are too small, and our networks are likely deriving an implicit advantage from being forcibly initialized to a convex function.
-Both methods are able to converge for learning rates in between $10^{-4}$ and $10^{-2}$; one could run for more epochs to see a similar advantage of our method for smaller learning rates, if desired.
+Both methods are able to converge for learning rates in between $10^{-4}$ and $10^{-2}$; one could run for more epochs to see a similar advantage of our method for smaller learning rates.
 All results in the main body of the paper used a constant learning rate of $10^{-3}$.
@@
 But this only accounts for losses on the order of $10^{-4}$, which indicates that in the more reasonable learning rate ranges, our pretraining is performing a meaningful function and enabling order-of-magnitude improvements over default network training.
-A.10 Real-World Data and Classification Problems
+B.2 Real-World Data and Classification Problems
 Here we present a few preliminary results on extending our pretraining technique to classification problems and real-world datasets.
@@
 In the case of classification, the loss function is simply swapped for cross entropy.
-After our pretraining phase, there are two choices of how to conduct the second phase of training (training matrix entries directly).
-All the parameters could be freed from constraints (“dense”), or the smaller 4-neuron subnetworks could be kept isolated from each other (but otherwise have their parameters freed).
-In the case where the 4-neuron networks remain in isolation, the weight matrices of the hidden layers will have a block diagonal structure (block size 4—the width of each subnetwork).
-Thus, we consider two fair (same number of free parameters) comparisons in this subsection: (1) dense versions of our and Kaiming-initialized networks, and (2) block diagonal versions of our and Kaiming-initialized networks.
-As discussed below and in the figure captions, we find that while the dense variant of our method ties or is slightly worse than a regular fully-connected network of the same dimensions, in the block diagonal case, our experimental networks significantly outperform their Kaiming-initialized counterparts.
-This makes sense in light of the experiments from the main body, where we can effectively shape the output of 4-neuron-wide networks better than random initializations, but where ensembles of our networks (see approximation of $x^{3}-x$ in Figure 6) get filled in with noise by gradient descent during the second training stage.
-This again highlights the need for more mathematical developments to provide a better extension into higher-dimensional nonconvex functions.
 Figure 15: Train/test loss plots for Kaiming-initialized (“ordinary”) fully-connected networks and our method (“experimental”).
@@
 We used 16 of our convex blocks, so all weight matrices are 64 by 64.
-Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 14 and 15 already show impressive results using our current approach.
+After our pretraining phase, there are two choices of how to conduct the second phase of training (training matrix entries directly).
+All the parameters could be freed from constraints (“dense”), or the smaller 4-neuron subnetworks could be kept isolated from each other (but otherwise have their parameters freed).
+In the case where the 4-neuron networks remain in isolation, the weight matrices of the hidden layers will have a block diagonal structure (block size 4—the width of each subnetwork).
+Thus, we consider two fair (same number of free parameters) comparisons in this subsection: (1) dense versions of our and Kaiming-initialized networks, and (2) block diagonal versions of our and Kaiming-initialized networks.
+As discussed below and in the figure captions, we find that while the dense variant of our method ties or is slightly worse than a regular fully-connected network of the same dimensions, in the block diagonal case, our experimental networks significantly outperform their Kaiming-initialized counterparts.
+This makes sense in light of the experiments from the main body, where we can effectively shape the output of 4-neuron-wide networks better than random initializations, but where ensembles of our networks (see approximation of $x^{3}-x$ in Figure 6) get filled in with noise by gradient descent during the second training stage.
+This again highlights the need for more mathematical developments to provide a better extension into higher-dimensional nonconvex functions.
+Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 15 and 16 already show impressive results using our current approach.
 In the case of the concrete problem (Figure 15), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced.
@@
 Further interesting optimizations of our approach (e.g., training another network to inform our algorithm when to switch parameterizations) are imagined as future work.
-A.11 Larger Image Classification Examples
-We include very preliminary demonstrations of how our method can increase performance on large-scale image classification tasks.
+B.3 Image Classification on CIFAR-10
+Figure 17: Results for CIFAR-10 experiments.
+Our (“experimental”) training and validations are generally comparable to their Kaiming-initialized (“ordinary”) counterparts.
+The small architectures are trained with Adam at a learning rate of $0.001$, while the VGG experiments are trained with SGD, following a learning rate schedule with a 10x drop at epochs 80 and 120.
+The ordinary network uses an initial learning rate of 0.1, and the experimental network uses 0.01.
+No parameterization switches occur with our method.
 A nice feature of our method is that it can be used anywhere dense layers appear in networks, which enables its use in CNN architectures.
-Nonetheless, combining our construction with convolutions or other layers is a use case far outside of the scope of some of the toy problems where we observe orders of magnitude of improvement, and the output of each subnetwork might not learn to be as nicely behaved as in the tightly controlled experimental setting of 1D convex training data.
-Thus, we stress that while our initial results are positive, and we can reasonably conjecture why, future work on extending our method to arbitrary ReLU networks would provide greater rigor and likely far greater performance gains, as seen with our lower-dimensional function approximation examples that realize the full benefit of our mathematically principled approach.
-We evaluated two standard benchmarks: CIFAR-10 and ImageNet.
-In the case of CIFAR-10, we trained a small network of three convolutions followed by two fully connected layers.
-To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added.
-In the case of ImageNet, we used PyTorch’s pretrained VGG16 model for its convolution weights (which were frozen during training).
-We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization.
-We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImageNet.
-Note that in both experiments, we switch parameterizations in our method halfway through training.
-Figure 16: results for CIFAR-10 and ImageNet experiments.
-Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts.
-Interestingly, the results on ImageNet are much stronger (note the color labeling is the same for both images, our train and validation losses are the bottom two in ImageNet, whereas in Cifar-10 the two training losses are the lowest).
-This may be because the ImageNet experiments re-used the convolutional layers from VGG, rather than learning them on its own.
-It could also be in part due to larger layer sizes, since most activation functions are universal with infinite width, any detrimental effect of constraining them into convex shapes may be softened.
-In our ImageNet experiment we trained both networks for 2 epochs.
-Our method required a lower learning rate ($0.00001$) than used in previous experiments.
-At epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).
-These two examples show that even for large-scale image classification tasks—far away from the analytic 1D function tests considered in the main body—our methodology still has the potential to yield noticeable benefits, highlighting the promise of extending the present work to more general network architectures as a future research direction.
-Potential Broader Impact
-This paper presents work whose goal is to enable more efficient neural networks.
-Although the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
+Here we demonstrate its application for image classification on CIFAR-10 (Krizhevsky et al., 2009).
+Our networks are inserted into the dense layers of VGG-16, as well as a small-scale CNN architecture of 3 convolutions (channel depths of $3\rightarrow 32\rightarrow 64\rightarrow 128$ with $3\times 3$ kernels) followed by a single hidden layer dense classifier of 256 neurons, whose width was reduced slightly to 247 to pay for the extra parameters when our method was used.
+The results we get are comparable between the two approaches (similar to the ImageNet experiments in Section 6).
+Again we notice that our method requires a lower learning rate for stability, and that switching parameterizations does not significantly impact accuracy.
+Appendix C Source Code Release
+An implementation of our technique can be found at https://github.com/NREL/triangle_net.